### v2

# Exact Subgraph Isomorphism Network with Mixed $L_{0,2}$ Norm Constraint for Predictive Graph Mining

Abstract
In the graph-level prediction task (predict a label for a given graph), the information contained in subgraphs of the input graph plays a key role. In this paper, we propose Exact subgraph Isomorphism Network (EIN), which combines the exact subgraph enumeration, a neural network, and a sparse regularization by the mixed $L_{0,2}$ norm constraint. In general, building a graph-level prediction model achieving high discriminative ability along with interpretability is still a challenging problem. Our combination of the subgraph enumeration and neural network contributes to high discriminative ability about the subgraph structure of the input graph. Further, the sparse regularization in EIN enables us 1) to derive an effective pruning strategy that mitigates computational difficulty of the enumeration while maintaining the prediction performance, and 2) to identify important subgraphs that contributes to high interpretability. We empirically show that EIN has sufficiently high prediction performance compared with standard graph neural network models, and also, we show examples of post-hoc analysis based on the selected subgraphs.
1 Introduction
Graph-level prediction tasks, which take a graph as an input and predict a label for the entire graph, have been widely studied in the data-science community. It is known that the graph representation is an effective approach to a variety of structure data such as chemical compounds (Ralaivola et al.,, 2005; Faber et al.,, 2017), protein structures (Gligorijević et al.,, 2021), and inorganic crystal structures (Xie and Grossman,, 2018; Louis et al.,, 2020). In a graph-level prediction task, substructures on the input graph, i.e., subgraphs, are often an important factor for the prediction and the analysis. For example, in a prediction of a property of chemical compounds, identifying small substructures of the molecules can be essential for both of improving prediction accuracy and obtaining an insight about the underlying chemical mechanism. Therefore, mining predictive subgraphs is a significant issue for graph-level prediction tasks, and further, those subgraphs can have a higher order dependency to the prediction that requests sufficient flexibility in the model. However, building a prediction method that satisfies these requirements is a still challenging problem (see § 3 for existing studies).
Our proposed method, called Exact subgraph Isomorphism Network (EIN), adaptively identifies predictive subgraphs based on which a neural network model can be simultaneously trained. The overview of EIN is shown in Fig. 1. Fig. 1 (a) illustrates the subgraph representation of EIN denoted as $\psi_{H}(G)$ which takes a non-zero value if the input graph $G$ contains a subgraph $H$ and takes 0 otherwise (i.e., it is based on the exact subgraph isomorphism). As shown in Fig. 1 (b), EIN can be seen as a neural network in which the candidates of the input features are all the subgraphs contained in the training dataset. Because of its exact subgraph isomorphism representation, this architecture can be highly discriminative about subgraph structures. However, since the number of the candidate subgraphs can be enormous, the naïve computation of this architecture is computationally intractable.
Figure 1: Overview of proposed method.
EIN consists of Graph Mining Layer (GML) and Feed Forward Network (FFN). GML implements a mechanism that identifies only $s$ subgraphs important for the prediction. This is achieved through a group sparse constraint (Yuan and Lin,, 2006), where $s$ is a pre-specified small integer (typically, less than $100$). We regard a set of neural network weights from an input unit (corresponding to one subgraph) to the next layer as a group, by which the adaptive subgraph selection can be realized. Only the selected subgraphs are used in the subsequent FFN that learns the dependency between the selected subgraphs and the prediction target. Although the number of subgraphs becomes small in the trained model, this approach still suffers from the computational difficulty because all the candidate subgraphs are still required to consider during the optimization process.
To tackle the computational difficulty, we combine a subgraph enumeration by graph mining (Yan and Han,, 2002) and the iterative hard thresholding (IHT) (e.g., Blumensath and Davies,, 2008; Herrity et al.,, 2006; Jain et al.,, 2014; Damadi et al.,, 2024), by which a pruning strategy for unnecessarily subgraphs can be derived, as illustrated in 1 (c). In IHT, only when the norm of the gradient corresponding to a subgraph $H$ ranks within the top $s$-th largest values among all the possible subgraphs, parameters for that subgraph is required to update. We show that, by deriving an upper bound of the norm of the gradient, an efficient pruning strategy for the subgraph enumeration can be constructed. Our pruning strategy has the following two important implications. First, this pruning enables us to train EIN without enumerating all the candidate subgraphs, which makes EIN computationally tractable. Second, our pruning strategy maintains the quality of the prediction compared with when we do not perform the pruning. This is because we only omit the computations that does not have any effect on the prediction.
Our contributions are summarized as follows.
  • We propose EIN, which is a neural network model that uses the exact subgraph isomorphism feature. Through a group sparse constraint, we formulate EIN so that a small number of subgraphs can be identified by which an insight about the important substructure can be extracted for the given graph-level prediction task.
  • We show that the combination of the subgraph enumeration and IHT with backpropagation can derive an efficient pruning strategy that makes the EIN training computationally tractable. We further reveal that our pruning strategy does not degrade the prediction quality, and the resulting algorithm has at least sublinear convergence rate to a critical point.
  • Based on synthetic and benchmark datasets, we demonstrate that EIN has superior or comparable performance compared with standard graph prediction models while EIN actually can identify a small number of important subgraphs, simultaneously.
2 Proposed Method: Exact Subgraph Isomorphism Network
Our proposed method, called Exact subgraph Isomorphism Network (EIN), considers the classification problem of a graph $G\in{\mathcal{G}}$, where $\mathcal{G}$} is a set of labeled graphs. $G$ consists of a set of nodes and edges between nodes and each node can have a categorical label. The training data is $\{(G_{i},y_{i})\}_{i\in[n]}$, where $y_{i}\in{\mathcal{Y}}$ is a graph label, $n$ is the number of instances, and $n]=\{1,\ldots,n\}$. Here, although we only focus on the classification problem, the regression problem can also be handled by just replacing the loss function.
First, § 2.1 describes the formulation of our model. Second, the optimization procedure is shown in § 2.2. Next, § 2.3 and 2.4 discuss post-hoc analysis for knowledge discovery and a combination of graph neural networks, respectively.
2.1 Model Definition
Let $\psi_{H}(G_{i})\in\{0,1\}$ be the feature that represents whether the input graph $G_{i}$ contains a subgraph $H$ (we only focus on a connected subgraph), which we call subgraph isomorphism feature (SIF):
$\psi_{H}(G_{i})=\mathbb{I}(H\sqsubseteq G_{i}),$ (1)
[…]
As shown in Fig. 1 (b), EIN consists of Graph Mining Layer (GML) and Feed Forward Network (FFN). Let $K$ be the number of output units of GML. We define $\mathrm{GML}:{\mathcal{G}}\rightarrow\mathbb{R}^{K}$ as follows.
$\mathrm{GML}(G_{i};\boldsymbol{B},\boldsymbol{b})$ $=\sigma(\boldsymbol{h}),$
$\boldsymbol{h}$ $=\sum_{H\in{\mathcal{H}}}\boldsymbol{\beta}_{H}\psi_{H}(G_{i})+\boldsymbol{b},$
where $\boldsymbol{B}\in\mathbb{R}^{K\times|{\mathcal{H}}|}$ is a weight matrix whose columns consist of $\boldsymbol{\beta}_{H}\in\mathbb{R}^{K}$ for $H\in{\mathcal{H}}$, $\boldsymbol{b}\in\mathbb{R}^{K}$ is a bias parameter, and $\sigma:\mathbb{R}^{K}\rightarrow\mathbb{R}^{K}$ is an activation function (this time, we use the standard sigmoid function). Each $\boldsymbol{\beta}_{H}$ can be seen as a representation corresponding to each subgraph $H$. The entire prediction model is defined as
$f(G_{i})$ $=\mathrm{FFN}(\mathrm{GML}(G_{i};\boldsymbol{B},\boldsymbol{b});\boldsymbol{\theta}),$
where $\boldsymbol{\theta}$ is parameters of FFN. We use the cross entropy loss function by which our objective function is defined as $\ell(\boldsymbol{B},\boldsymbol{b},\boldsymbol{\theta})\coloneqq\sum_{i=1}^{n}\mathrm{loss}(y_{i},f(G_{i});\boldsymbol{B},\boldsymbol{b},\boldsymbol{\theta})$. We optimize the parameters $\boldsymbol{B},\boldsymbol{b}$, and $\boldsymbol{\theta}$ through the following optimization problem in which a group sparsity constraint is imposed on $\boldsymbol{B}$:
$\begin{split}\min_{\boldsymbol{B},\boldsymbol{b},\boldsymbol{\theta}}\ &\ell(\boldsymbol{B},\boldsymbol{b},\boldsymbol{\theta})\\ \text{s.t.}\ &\|\boldsymbol{B}\|_{0,2}\leq s,\end{split}$ (2)
where
$\|\boldsymbol{B}\|_{0,2}=|\{H\mid\|\boldsymbol{\beta}_{H}\|_{2}\neq 0\}|$
is the mixed $L_{0,2}$ norm, and $s>0$ is the constant parameter. The constraint in (2) indicates that $\boldsymbol{B}$ only has $s$ non-zero columns. In other words, this means that only $s$ subgraphs are used in $f$, by which we can identify important predictive subgraphs as $\{H\mid\boldsymbol{\beta}_{H}\neq\boldsymbol{0}\}$. However, since the size of $\boldsymbol{B}$, i.e., $K\times|{\mathcal{H}}|$, is quite large, naïve optimization of (2) can be difficult.
2.2 Optimization
Our optimization algorithm is based on the block coordinate descent, in which each one of $\boldsymbol{B},\boldsymbol{b}$, and $\boldsymbol{\theta}$ are updated alternately while the other two parameters are fixed. Since $\boldsymbol{B}$ has the group sparse constraint, we apply iterative hard thresholding (IHT) to $\boldsymbol{B}$ (e.g., Beck and Eldar,, 2013; Bao et al.,, 2014; Yuan et al.,, 2018; Damadi and Shen,, 2022). On the other hand, $\boldsymbol{b}$ and $\boldsymbol{\theta}$ have no sparse penalty and we can simply apply the standard gradient descent.
We iterate the following procedure, in which $t$ is the number of iterations:
  1. 1.
    Update $\boldsymbol{B}$ through the hard thresholding (HT), defined as
    $\begin{split}\boldsymbol{B}^{t+\frac{1}{2}}&=\boldsymbol{B}^{t}-\gamma^{t}_{\boldsymbol{B}}\nabla_{\boldsymbol{B}}\ell^{t}\\ \boldsymbol{B}^{t+1}&=H_{s}(\boldsymbol{B}^{t+\frac{1}{2}}),\end{split}$ (3)
    where $\nabla_{\boldsymbol{B}}\ell^{t}={\frac{\partial\ell(\boldsymbol{B}^{t},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})}{\partial\boldsymbol{B}}}$ is the gradient of $\boldsymbol{B}$, $\gamma^{t}_{\boldsymbol{B}}$ is a step length, and
    $H_{s}(\boldsymbol{B})=\mathop{\text{argmin}}_{\|\boldsymbol{Z}\|_{0,2}\leq s}\|\boldsymbol{Z}-\boldsymbol{B}\|_{F}$
    is the HT operator. The step length parameter $\gamma^{t}_{\boldsymbol{B}}$ is determined by the backtrack search (Nocedal and Wright,, 2006).
  2. 2.
    Update $\boldsymbol{b}^{t}$ and $\boldsymbol{\theta}^{t}$ to $\boldsymbol{b}^{t+1}$ and $\boldsymbol{\theta}^{t+1}$, respectively, by using the standard gradient descent.
Since $\mathcal{H}$} contains a large number of subgraphs, calculating (3) for the entire $\boldsymbol{B}\in\mathbb{R}^{K\times|{\mathcal{H}}|}$ is not directly tractable. In § 2.2.1, we derive a pruning rule by which we can perform (3) without enumerating all the subgraphs in $\mathcal{H}$}. § 2.2.2 describes the entire procedure, and § 2.2.3 provides a theoretical convergence analysis.
2.2.1 Backpropagation with Gradient Pruning
By combining a graph mining algorithm and the HT operation, we can build an efficient pruning algorithm. We employ a well-known graph mining algorithm called gSpan (Yan and Han,, 2002) that can efficiently enumerate all the subgraphs in a given set of graphs. Figure 1 (c) is an illustration of a graph mining tree. At each tree node, gSpan expands the graph (add an edge and a node) as far as the expanded graph is included in the given dataset as a subgraph. By providing a unique code to each generated graphs (called the DFS code because it is based on depth-first search in a graph), gSpan can enumerate all the subgraphs without generating duplicated graphs (For further detail, see (Yan and Han,, 2002)). As shown in Fig. 1 (c), all the graphs in the tree contains their ancestors as a subgraph.
It is easy to see that the HT operator simply selects the columns with the top-$s$ largest norm because
$H_{s}(\boldsymbol{B})$ $=\mathop{\text{argmin}}_{\|\boldsymbol{Z}\|_{2,0}\leq s}\|\boldsymbol{Z}-\boldsymbol{B}\|_{F}^{2}=\mathop{\text{argmin}}_{\|\boldsymbol{Z}\|_{2,0}\leq s}\sum_{k=1}^{|{\mathcal{H}}|}\|\boldsymbol{Z}_{k}-\boldsymbol{B}_{k}\|_{2}^{2},$
where $\boldsymbol{B}_{k}$ and $\boldsymbol{Z}_{k}$ is the $k$-th column of $\boldsymbol{B}$ and $\boldsymbol{Z}$, respectively. Let $H_{1},\ldots,H_{|{\mathcal{H}}|}$ be the subgraphs such that $\|\boldsymbol{\beta}^{t+\frac{1}{2}}_{H_{1}}\|_{2}>\|\boldsymbol{\beta}^{t+\frac{1}{2}}_{H_{2}}\|_{2}>\cdots>\|\boldsymbol{\beta}^{t+\frac{1}{2}}_{H_{|{\mathcal{H}}|}}\|_{2}$, where $\boldsymbol{\beta}^{t+\frac{1}{2}}_{H}$ is a column of $\boldsymbol{B}^{t+\frac{1}{2}}$. Then, $\boldsymbol{\beta}^{t+1}_{H}$, which is a column of $\boldsymbol{B}^{t+1}$, is written as
$\boldsymbol{\beta}^{t+1}_{H}=\begin{cases}\boldsymbol{\beta}^{t+\frac{1}{2}}_{H}&\text{ if }H\in\{H_{1},\ldots,H_{s}\},\\ \boldsymbol{0}&\text{ otherwise. }\end{cases}$
This means that we only need to identify $H$ with the top-$s$ largest $\|\boldsymbol{\beta}^{t+\frac{1}{2}}_{H}\|_{2}$ during the graph mining tree traverse because $\boldsymbol{\beta}^{t+1}_{H}$ for the other $H$ is $\boldsymbol{0}$.
Let $S_{t}$ be the set of selected $H$ at the iteration $t$. Since $|S_{t}|=s$, which is assumed to be a small value (in our later experiments, $s\leq 100$), it is not computationally difficult to calculate $\boldsymbol{\beta}^{t+\frac{1}{2}}_{H}$ for $H\in S_{t}$. On the other hand, we have an intractable number of $H\notin S_{t}$. Therefore, we remove $H$ that does not have a possibility of being included in $S_{t+1}$. Since $\boldsymbol{\beta}^{t}_{H}=\boldsymbol{0}$ for $H\notin S_{t}$, if $H\notin S_{t}$, we see $\|\boldsymbol{\beta}_{H}^{t+\frac{1}{2}}\|_{2}=\gamma_{\boldsymbol{B}}^{t}\|\nabla_{\boldsymbol{\beta}_{H}}\ell^{t}\|_{2}$, where $\nabla_{\boldsymbol{\beta}_{H}}\ell^{t}={\frac{\partial\ell(\boldsymbol{B}^{t},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})}{\partial\boldsymbol{\beta}_{H}}}$. To define our pruning strategy, the following upper bound of $\|\nabla_{\boldsymbol{\beta}_{H}}\ell^{t}\|_{2}$ plays a key role:
Theorem 2.1.
Let $H^{\prime}\sqsupseteq H$ and $\delta^{t}_{ik}={\frac{\partial\mathrm{loss}(y_{i},f(G_{i});\boldsymbol{B}^{t},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})}{\partial h_{k}}}$. Then,
$\|\nabla_{\boldsymbol{\beta}_{H^{\prime}}}\ell^{t}\|_{2}\leq$ $\biggl\{\sum_{k=1}^{K}\max\bigl\{\bigl(\sum_{i\in\{i\mid\delta^{t}_{ik}>0\}}\delta^{t}_{ik}\psi_{H}(G_{i})\bigr)^{2},$
$\ \ \bigl(\sum_{i\in\{i\mid\delta^{t}_{ik}<0\}}\delta^{t}_{ik}\psi_{H}(G_{i})\bigr)^{2}\bigr\}\biggr\}^{1/2}\eqqcolon\mathrm{UB}(H).$ (4)
We here only describe the sketch of the proof (the proof is in Appendix C). An essential idea is that we expand $\nabla_{\boldsymbol{\beta}_{H}}\ell^{t}$ as a linear combination of the derivatives with respect to the GML intermediate representation $h_{k}$, i.e., $\delta^{t}_{ik}$, by which the $k$-th element of the gradient can be written as $(\nabla_{\boldsymbol{\beta}_{H}}\ell^{t})_{k}=\sum_{i=1}^{n}\delta^{t}_{ik}\psi_{H}(G_{i})$. By using the monotonically non-increasing property of $\psi_{H}(G_{i})$, i.e., $\psi_{H^{\prime}}(G_{i})\leq\psi_{H}(G_{i})$ if $H\sqsubseteq H^{\prime}$, we can derive ${\rm UB}(H)$. Note that, usually, there is no need to explicitly consider this specific expansion of $(\nabla_{\boldsymbol{\beta}_{H}}\ell^{t})_{k}$; the theorem can only be derived by deliberately reducing it to the linear combination of $\delta^{t}_{ik}$.
Theorem 2.1 indicates that, for any $H^{\prime}$ that contains $H$ as a subgraph, the $L_{2}$ norm of the gradient $\|\nabla_{\boldsymbol{\beta}_{H^{\prime}}}\ell^{t}\|_{2}$ can be bounded by $\mathrm{UB}(H)$. Note that the upper bound $\mathrm{UB}(H)$ can be calculated without generating $H^{\prime}$. From Theorem 2.1, we can immediately obtain the following pruning rule with a threshold parameter $\eta$:
Corollary 2.1.
Let $\zeta\leq\|\boldsymbol{\beta}^{t+\frac{1}{2}}_{H_{s}}\|_{2}/\gamma_{\boldsymbol{B}}^{t}$, meaning that $\zeta$ is less than or equal to the $s$-th largest value of $\|\boldsymbol{\beta}^{t+\frac{1}{2}}_{H}\|_{2}/\gamma_{\boldsymbol{B}}^{t}$ among $H\in{\mathcal{H}}$. For $\forall H^{\prime}\in\{H^{\prime}\mid H^{\prime}\sqsupseteq H,\ H^{\prime}\notin S_{t}\}$,
$\mathrm{UB}(H)<\zeta\ \ \Rightarrow\ \ \boldsymbol{\beta}_{H^{\prime}}^{t+1}=\boldsymbol{0}$ (5)
This is a direct consequence from the fact that if the conditions in (5) holds, we have
$\|\boldsymbol{\beta}^{t+\frac{1}{2}}_{H^{\prime}}\|_{2}=\gamma_{\boldsymbol{B}}^{t}\|\nabla_{\boldsymbol{\beta}_{H^{\prime}}}\ell^{t}\|_{2}\leq\gamma_{\boldsymbol{B}}^{t}\mathrm{UB}(H)<\gamma_{\boldsymbol{B}}^{t}\zeta\leq\|\boldsymbol{\beta}^{t+\frac{1}{2}}_{H_{s}}\|_{2}$
from (4). As a result, we see that $\|\boldsymbol{\beta}^{t+\frac{1}{2}}_{H^{\prime}}\|_{2}$ cannot be ranked in the top-$s$ largest values. Therefore, we can omit the update for $\forall H^{\prime}$ that contains $H$ as a subgraph if $\mathrm{UB}(H)<\zeta$ holds.
To use the pruning rule Corollary 2.1, we need to determine $\zeta$. We first consider the case that the step length $\gamma_{\boldsymbol{B}}^{t}$ is fixed beforehand, in which a simple approach can be applied. First, $\|\boldsymbol{\beta}_{H}^{t+\frac{1}{2}}\|$ is calculated for all $H\in S_{t}$. After that, the graph mining tree is traversed during which each one of $\|\boldsymbol{\beta}_{H}^{t+\frac{1}{2}}\|$ is calculated for the subgraph $H$ of each visited tree node. Suppose that ${\mathcal{H}}_{\rm visited}$ represents the union of $S_{t}$ and a set of $H$ corresponding to the already visited tree nodes. The threshold of the pruning rule (5) can be set as $\zeta=s\operatorname{-max}\{\|\boldsymbol{\beta}_{H}^{t+\frac{1}{2}}\|/\gamma_{\boldsymbol{B}}^{t}\}_{H\in{\mathcal{H}}_{\rm visited}}$, where $s\operatorname{-max}$ represents the $s$-th largest value, by which we can guarantee $\zeta\leq\|\boldsymbol{\beta}^{t+\frac{1}{2}}_{H_{s}}\|_{2}/\gamma_{\boldsymbol{B}}^{t}$. As a result, if $\mathrm{UB}(H)<\zeta$ holds for the threshold $\zeta$ at any node during the tree traverse, we can discard the entire subtree below the current node. However, when we use a backtrack strategy for the step length, we cannot calculate $\|\boldsymbol{\beta}_{H}^{t+\frac{1}{2}}\|_{2}$ before the update of $\boldsymbol{B}^{t}$ because $\gamma_{\boldsymbol{B}}^{t}$ is not determined yet.
Instead, we can also build a threshold $\zeta$ without using $\gamma_{\boldsymbol{B}}^{t}$ as follows.
$s\operatorname{-max}\{\ \|\boldsymbol{\beta}_{H}^{t+\frac{1}{2}}\|_{2}/\gamma_{\boldsymbol{B}}^{t}\ \}_{H\in{\mathcal{H}}_{\rm visited}}$
$=s\operatorname{-max}\left\{\{\frac{1}{\gamma_{\boldsymbol{B}}^{t}}\|\boldsymbol{\beta}_{H}^{t}-\gamma_{\boldsymbol{B}}^{t}\nabla_{\boldsymbol{\beta}_{H}}\ell^{t}\|_{2}\}_{H\in S_{t}}\cup\{\|\nabla_{\boldsymbol{\beta}_{H}}\ell^{t}\|_{2}\}_{H\in{\mathcal{H}}_{\rm visited}\setminus S_{t}}\right\}$
$\geq s\operatorname{-max}\left\{\eta_{H}\right\}_{H\in{\mathcal{H}}_{\rm visited}},$ (6)
where
$\eta_{H}\coloneqq\begin{cases}\min_{\gamma\in\Gamma}\frac{1}{\gamma}\|\boldsymbol{\beta}_{H}^{t}-\gamma\nabla_{\boldsymbol{\beta}_{H}}\ell^{t}\|_{2}&\text{ if }H\in S_{t},\\ \|\nabla_{\boldsymbol{\beta}_{H}}\ell^{t}\|_{2}&\text{ if }H\notin S_{t},\end{cases}$
and $\Gamma$ is a candidate space of the step length $\gamma_{\boldsymbol{B}}^{t}$. We use this lower bound (6) as $\zeta$. In the derivation of (6), $H\in S_{t}$ and other $H\in{\mathcal{H}}_{\rm visited}\setminus S_{t}$ are separately considered. For $H\in{\mathcal{H}}_{\rm visited}\setminus S_{t}$ (i.e., $H\notin S_{t}$), $\gamma_{\boldsymbol{B}}^{t}$ is vanished in $\eta_{H}$, and therefore, the calculation does not depend on the step length. On the other hand, for $H\in S_{t}$, since the step length cannot be directly removed, we derive a lower bound with respect to possible step lengths. This lower bound $\min_{\gamma\in\Gamma}$ can be easily calculated (See Appendix D for detail).
1function Train-EIN($\mathcal{S}$})
2 $\boldsymbol{B}\leftarrow\boldsymbol{0}$
3   Initialize $\boldsymbol{b}$ and $\boldsymbol{\theta}$ randomly
4 $H_{0}\leftarrow\text{a graph at root of mining tree}$
5 for $s$ in $\mathcal{S}$} do
6    repeat
7         Calculate $\eta_{H}$ for $\forall H\in S_{t}$
8       $S_{t+\frac{1}{2}}\leftarrow S_{t}$
9       $\zeta\leftarrow s\operatorname{-max}\left\{\eta_{H}\right\}_{H\in S_{t+\frac{1}{2}}}$
10       $S_{t+\frac{1}{2}}\leftarrow\mathrm{Traverse}(H_{0},s,\zeta,S_{t+\frac{1}{2}})$
11       $\boldsymbol{B}^{t+1}_{S_{t+\frac{1}{2}}}\leftarrow H_{s}(\boldsymbol{B}_{S_{t+\frac{1}{2}}}^{t}-\gamma_{\boldsymbol{B}}^{t}\nabla_{\boldsymbol{B}_{S_{t+\frac{1}{2}}}}\ell^{t})$, where $\gamma_{\boldsymbol{B}}^{t}$ is determined by backtrack search
12       $\boldsymbol{b}^{t+1}\leftarrow\mathrm{GradientDescent}(\boldsymbol{b}^{t})$
13       $\boldsymbol{\theta}^{t+1}\leftarrow\mathrm{GradientDescent}(\boldsymbol{\theta}^{t})$
14    until terminate condition met
15   end for
16
17 end function
18
Algorithm 1 Optimization for EIN
1function Traverse($H,s,\zeta,S_{t+\frac{1}{2}}$)
2   Calculate $\eta_{H}$ and $\mathrm{UB}(H)$ for current $H$
3 if $\mathrm{UB}(H)<\zeta$ then
    return $S_{t+\frac{1}{2}}$
[…]
4
5 else if $\eta_{H}>\zeta$ then
6    $S_{t+\frac{1}{2}}\leftarrow S_{t+\frac{1}{2}}\cup\{H\}$
7    $\zeta=s\operatorname{-max}\left\{\eta_{H^{\prime}}\right\}_{H^{\prime}\in S_{t+\frac{1}{2}}}$
8    $S_{t+\frac{1}{2}}\leftarrow S_{t+\frac{1}{2}}\setminus\{H^{\prime}\mid\eta_{H^{\prime}}<\zeta,\ H^{\prime}\notin S_{t}\}$
9   end if
10 ${\mathcal{C}}\leftarrow\mathrm{Expand}(H)$
11 for $H^{\prime}\in{\mathcal{C}}$ do
12    $S_{t+\frac{1}{2}}\leftarrow\mathrm{Traverse}(H,s,\zeta,S_{t+\frac{1}{2}})$
13   end for
14 return $S_{t+\frac{1}{2}}$
15 end function
16
17function Expand($H$)
18 if children of $H$ have never been created by gSpan then
19    $\mathcal{C}$$\leftarrow$ all graphs expanded from $H$ by gSpan
20      Set $\{\psi_{H^{\prime}}(G_{i})\}_{i=1}^{n}$ for $H^{\prime}\in{\mathcal{C}}$
21
22 else
23    $\mathcal{C}$$\leftarrow$ Retrieve already expanded children of $H$
24
25   end if
26 return $\mathcal{C}$}
27 end function
28
Algorithm 2 Traverse Mining Tree
As a result, the following important consequence is obtained
Remark 2.1.
By evaluating $\mathrm{UB}(H)\leq\zeta$ during the tree traverse of gSpan (e.g., depth-first search), all the descendants of $H$ in the tree can be pruned if the inequality holds. This means that we can perform the update (3) without enumerating all the possible subgraphs $\mathcal{H}$}.
Further, another notable remark about the pruning is as follows.
Remark 2.2.
Since our pruning strategy only omit the update of unnecessarily parameters, i.e., $\boldsymbol{\beta}_{H}=\boldsymbol{0}$, the pruning does no change any parameter values compared with the case that we do not use the pruning. This also means that the final prediction performance also does not change because of the pruning.
2.2.2 Algorithm
We here describe the optimization procedure of EIN. The basic strategy is that we first set $s$ as a small value (e.g., $1$) and gradually increase $s$ while optimizing parameters. Let ${\mathcal{S}}=(s_{1},\ldots,s_{K})$ be a sequence of $s$, where $s_{1}<s_{2}<\cdots<s_{K}$. For each $s$, we use the solution of previous $s$ as an initial solution.
The Train-EIN function of Algorithm 1 performs the optimization (2) for each $s$. In line 10, the Traverse function performs the graph mining tree search that generates a set of candidate subgraphs denoted as $S_{t+\frac{1}{2}}$, consisting of subgraphs that are not pruned (i.e., $S_{t+1}\subseteq S_{t+\frac{1}{2}}$). The detailed procedure of the candidate generation is in Algorithm 2. The Traverse function recursively searches the graph mining tree. At each tree node, $\mathrm{UB}(H)$ is evaluated and the entire subtree below the current node can be pruned if $\mathrm{UB}(H)<\zeta$ (line 3-4). On the other hand, if $\eta_{H}>\zeta$, $H$ is added to $S_{t+\frac{1}{2}}$ (line 6). Then, the threshold $\zeta$ is updated (line 7), after which we can remove $\{H^{\prime}\mid\eta_{H^{\prime}}<\zeta,\ H^{\prime}\notin S_{t}\}$ (line 8). This guarantees $|S_{t+\frac{1}{2}}|\leq 2s$ (for simplicity, we assume there is no tie of $\eta_{H}$). The children nodes of $H$ are created by gSpan in the Expand function, by which only the subgraphs $H^{\prime}$ included in the training dataset can be generated and as a byproduct of this process, we obtain $\psi_{H^{\prime}}(G_{i})$. Note that if the children are already generated in the previous iterations, we can reuse them. Once $S_{t+\frac{1}{2}}$ is determined, $\boldsymbol{B}^{t}$ is updated in line 11 of Algorithm 1, in which the step length $\gamma_{\boldsymbol{B}}^{t}$ is determined by the batcktrack search. Here, $\boldsymbol{B}_{S_{t+\frac{1}{2}}}$ is a submatrix of $\boldsymbol{B}$ in which only the columns specified by $S_{t+\frac{1}{2}}$ are extracted. In line 12-13 of Algorithm 1, $\boldsymbol{b}$ and $\boldsymbol{\theta}$ are updated by the standard gradient descent, in which the batcktrack search is also performed.
2.2.3 Convergence Analysis
We here show that EIN can converge to a critical point with at least sublinear convergence.
Suppose that the gradient descent of $\boldsymbol{b}$ and $\boldsymbol{\theta}$ iterates $\tau_{\max}$ times. The index $\tau$ represents the number of iteration of this inner loop as follows.
$\boldsymbol{b}^{t,\tau+1}$ $=\boldsymbol{b}^{t,\tau}-\gamma^{t,\tau}_{\boldsymbol{b}}\nabla_{\boldsymbol{b}}\ell^{t,\tau},$
$\boldsymbol{\theta}^{t,\tau+1}$ $=\boldsymbol{\theta}^{t,\tau}-\gamma^{t,\tau}_{\boldsymbol{\theta}}\nabla_{\boldsymbol{\theta}}\ell^{t,\tau},$
where $\nabla_{\boldsymbol{b}}\ell^{t,\tau}={\frac{\partial\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t,\tau},\boldsymbol{\theta}^{t})}{\partial\boldsymbol{b}}}$, $\nabla_{\boldsymbol{\theta}}\ell^{t,\tau}={\frac{\partial\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t+1},\boldsymbol{\theta}^{t,\tau})}{\partial\boldsymbol{\theta}}}$, and $\gamma^{t,\tau}_{\boldsymbol{b}}$ and $\gamma^{t,\tau}_{\boldsymbol{\theta}}$ are the step lengths of the $\tau$-th iteration, respectively. Let $c_{\boldsymbol{B}},c_{\boldsymbol{b}},c_{\boldsymbol{\theta}}\in(0,1)$ be constants used in the backtrack termination conditions. Our backtrack, in which the step length iteratively decays with a ratio $\rho\in(0,1)$, stops when
$\ell(\boldsymbol{B}^{t},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})-\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})$ $\geq{c_{\boldsymbol{B}}(1-\rho)\frac{\gamma^{t}_{\boldsymbol{B}}}{2}\|\nabla_{\boldsymbol{B}}\ell^{t}_{S}\|_{F}^{2}}$ (7)
$\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t,\tau},\boldsymbol{\theta}^{t})-\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t,\tau+1},\boldsymbol{\theta}^{t})$ $\geq c_{\boldsymbol{b}}\frac{\gamma^{t,\tau}_{\boldsymbol{b}}}{2}\|\nabla_{\boldsymbol{b}}\ell^{t,\tau}\|_{2}^{2}$ (8)
$\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t+1},\boldsymbol{\theta}^{t,\tau})-\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t+1},\boldsymbol{\theta}^{t,\tau+1})$ $\geq c_{\boldsymbol{\theta}}\frac{\gamma^{t,\tau}_{\boldsymbol{\theta}}}{2}\|\nabla_{\boldsymbol{\theta}}\ell^{t,\tau}\|_{2}^{2}$ (9)
are satisfied, respectively. These conditions are from the well-known Armijo condition (Nocedal and Wright,, 2006).
Throughout the proof, we assume differentiability and $L$-smoothness of $\ell$:
Definition 2.1.
Let $L$ be the Lipschitz constant and $\boldsymbol{z}$ be the vector that concatenates all the elements of $\boldsymbol{B}$, $\boldsymbol{b}$, and $\boldsymbol{\theta}$. The (restricted) $L$-smoothness of $\ell$ is defined as
$\left\|{\frac{\partial\ell(\boldsymbol{B},\boldsymbol{b},\boldsymbol{\theta})}{\partial\boldsymbol{z}}}-{\frac{\partial\ell(\boldsymbol{B}^{\prime},\boldsymbol{b}^{\prime},\boldsymbol{\theta}^{\prime})}{\partial\boldsymbol{z}}}\right\|_{2}\leq L\|\boldsymbol{z}-\boldsymbol{z}^{\prime}\|_{2}$ (10)
for any pair of $(\boldsymbol{B},\boldsymbol{b},\boldsymbol{\theta})$ and $(\boldsymbol{B}^{\prime},\boldsymbol{b}^{\prime},\boldsymbol{\theta}^{\prime})$, where $\|\boldsymbol{B}\|_{0}\leq s$, $\|\boldsymbol{B}^{\prime}\|_{0}\leq s$.
We first guarantee the existence of the positive step lengths that satisfy the backtrack conditions.
Lemma 2.1.
There exist positive step lengths $\gamma_{\boldsymbol{B}}^{t}\geq\rho^{2}/L$, $\gamma_{\boldsymbol{b}}^{t,\tau}\geq\rho/L$, and $\gamma_{\boldsymbol{\theta}}^{t,\tau}\geq\rho/L$ that satisfy (7), (8), and (9), respectively.
The proof is in Appendix E. For $\boldsymbol{B}$, the proof is mainly from (Damadi and Shen,, 2022), and for $\boldsymbol{b}$ and $\boldsymbol{\theta}$, it is from a known property of the gradient descent. The convergence of our procedure is shown by the following theorem, in which our problem (2) is re-written as an equivalent unconstrained problem:
Theorem 2.2.
Let
$\tilde{\ell}(\boldsymbol{z})=\ell(\boldsymbol{B},\boldsymbol{b},\boldsymbol{\theta})+\delta_{s}(\boldsymbol{B})$
be the objective function with the penalty on the constraint violation
$\delta_{s}(\boldsymbol{B})=\begin{cases}0,&\text{ if }\|\boldsymbol{B}\|_{2,0}\leq s,\\ \infty,&\text{ otherwise. }\end{cases}$
The sequence $\boldsymbol{z}^{t}$, defined by $\boldsymbol{B}^{t}$, $\boldsymbol{b}^{t}$, and $\boldsymbol{\theta}^{t}$, is assumed to be bounded. Then, $\boldsymbol{z}^{t}$ is a Cauchy sequence that converges to a critical point $\boldsymbol{z}^{*}$ satisfying $\partial\tilde{\ell}(\boldsymbol{z}^{*})\ni\boldsymbol{0}$, where $\partial\tilde{\ell}(\boldsymbol{z})$ is the limiting subdifferential of $\tilde{\ell}$.
See Appendix G for the proof. The limiting subdifferential (Mordukhovich,, 2012) is a generalized notion of subdifferential applicable to an arbitrary proper and lower semi-continuous (extended-real-valued) function (See Appendix F for the definition). The generalized Fermat’s rule (Rockafellar et al.,, 2009) indicates $\partial\tilde{\ell}(\boldsymbol{z}^{*})\ni\boldsymbol{0}$ is a necessarily condition of a local minimum. The boundedness assumption of $\boldsymbol{z}^{t}$ is introduced to ensure that $\boldsymbol{z}^{t}$ does not diverge to infinity. In practice, if one wishes to guarantee this rigorously, a possible approach is to introduce an $L_{2}$ penalty (weight decay) on the parameters (for which the regularization coefficient can be chosen to be arbitrarily small). We further derive the following convergence rate:
Theorem 2.3.
Under the same assumption as Theorem 2.2, our optimization algorithm has at least the following sublinear convergence rate:
$\|\boldsymbol{z}^{t}-\boldsymbol{z}^{*}\|_{2}=O(t^{q}),$
where $q<0$.
Based on properties of $\boldsymbol{z}^{t}$ obtained through the proof of Theorem 2.2, the above rate can be derived by the almost same proof as Theorem 2 of (Attouch and Bolte,, 2009). We show the proof in Appendix H for completeness.
2.3 Post-hoc Analysis for Knowledge Discovery
After the optimization, only $s$ subgraphs have non-zero values of $\boldsymbol{\beta}_{H}$, by which we can identify predictive subgraphs. In our later experiments, we observed that small $s$ (up to $100$) can achieve sufficiently high predictive performance. Therefore, examining the selected subgraphs—possibly by a domain expert—may lead to new insights about the data.
The trained EIN can be seen as an $s$-dimensional input neural network, to which post-hoc knowledge discovery methods (Bhati et al.,, 2025) can be applied. For example, well-known SHAP (Lundberg and Lee,, 2017) and LIME (Ribeiro et al.,, 2016), both of which provide local feature importance, is applicable. Another typical approach is to use the trained EIN as a teacher and fit an interpretable surrogate model (e.g., decision tree), from which a possible decision rule can be estimated (Molnar,, 2025). Note that, usually, applying interpretable machine learning methods directly to exhaustive subgraph features can be computationally intractable. Benefiting from EIN, we can only focus on a small number of important selected subgraphs, by which a variety of post-hoc analyses become much easier to apply.
Figure 2: A simple example of EIN combined with GNN.
2.4 Combining with Graph Neural Network
Since EIN is based on the standard backpropagation mechanism, general neural network models can be combined flexibly. For example, a simple approach to combining EIN with a graph neural network (GNN) is
$\begin{split}f(G_{i})&=\mathrm{FFN}(\mathrm{GNN}(G_{i};\boldsymbol{\theta}_{\rm GNN})\\ &\qquad\oplus\mathrm{GML}(G_{i};\boldsymbol{B},\boldsymbol{b});\boldsymbol{\theta}_{\rm FFN}),\end{split}$ (11)
where $\oplus$ is the vector concatenation, $\mathrm{GNN}$ is any GNN that outputs arbitrary dimension representation vector, and $\boldsymbol{\theta}_{\rm GNN}$ and $\boldsymbol{\theta}_{\rm FFN}$ are parameters of $\mathrm{GNN}$ and $\mathrm{FFN}$, respectively. This combination enables any GNNs to enhance the discriminative ability in terms of the selected subgraphs by EIN. A schematic illustration of a simplest case is shown in Fig. 2. Computations of the combined model can also be performed by almost the same alternating update shown in Algorithm 1. The only difference is that the FFN parameter update (line 13 of Algorithm 1) is replaced with the simultaneous update of $\boldsymbol{\theta}_{\rm FFN}$ and $\boldsymbol{\theta}_{\rm GNN}$.
3 Related Work
For GNNs, the message passing based approach has been widely employed (Zhou et al.,, 2020). It is known that the expressive power of classical message passing approaches is limited by the first order Weisfeiler-Lehman (1-WL) test (Zhao et al.,, 2022). Many studies have tackled this limitation, some of which actually can be comparable with higher order WL tests. For example, $k$-hop extensions of the message passing (e.g., Abu-El-Haija et al.,, 2019; Nikolentzos et al.,, 2020; Wang et al.,, 2021; Chien et al.,, 2021; Brossard et al.,, 2020) and the higher order (tensor) representations (e.g., Morris et al.,, 2019; Maron et al., 2019b, ; Maron et al., 2019a, ; Maron et al., 2019c, ; Keriven and Peyré,, 2019; Geerts and Reutter,, 2022; Murphy et al.,, 2019) are popular approaches. Further, several studies (Bouritsas et al.,, 2023; Cotta et al.,, 2021; Barcelo et al.,, 2021; Bevilacqua et al.,, 2022; Yan et al.,, 2024) incorporate small subgraph information, such as pre-specified motifs and ego-networks, into neural networks. On the other hand, EIN performs the data-adaptive selection of a small number of (globally) predictive subgraphs, based on the exact enumeration of all the subgraphs in the dataset, which is obviously different approach from the above popular GNN studies. Further, most of GNNs can be combined with EIN by (11).
Although explainable GNNs are also studied, according to (Yuan et al.,, 2022), most of them are for the ‘instance-level’ explanation, which considers an explanation for the prediction of the specific one input graph. A few exceptions are approaches based on the maximization of the trained GNN output for a target label (Yuan et al.,, 2020), the latent space prototype learning (Azzolin et al.,, 2023), and the kernel-based filtering (Feng et al.,, 2022). However, unlike EIN, they cannot guarantee that the identified important graphs are actually subgraphs of the input graphs because none of them are based on the exact matching.
In the context of graph mining, discriminative pattern mining has also been studied (e.g., Thoma et al.,, 2010; Potin et al.,, 2025; Chen et al.,, 2022). However, most of them are not directly incorporated in a learning model. Typically, the subgraph enumeration (using some pruning strategies) is performed based on a discriminative score (e.g., ratio of frequency between two classes) independent from a learning model, and the selection results can be used in a subsequent prediction model training. Although (Nakagawa et al.,, 2016; Yoshida et al.,, 2019, 2021; Tajima et al.,, 2024) enumerate subgraphs during a model optimization, they are limited to a linear combination of the subgraphs. Although (Nakagawa et al.,, 2016; Yoshida et al.,, 2019, 2021) provide a stronger pruning rule that can satisfy the global optimality, the convex formulation of the model training is required. Several studies have considered learning $L_{0}$ based sparse neural networks (e.g., Louizos et al.,, 2018; Damadi et al.,, 2024). However, to our knowledge, a discriminative subgraph mining that directly combines the exact subgraph enumeration with an $L_{0,2}$ constraint neural network has not been investigated.
4 Experiments
We verify the prediction performance and interpretability of EIN through synthetic and benchmark datasets. For all the datasets, we partitioned them into ${\rm train}:{\rm valid}:{\rm test}=6:2:2$.
Synthetic Datasets.
(a) $H_{p}$
(b) $H_{n}$
(c) $H_{\rm padding}$
Figure 3: Subgraphs (a) and (b) have closed paths whose lengths are 8 and 9, respectively, which are difficult to discriminate. Subgraph (c) is used for adjusting the number of nodes that have label ‘1’. Note that all subgraphs (a), (b), and (c) have 16 nodes.
We created two synthetic datasets by using subgraphs shown in Figure 3. $H_{p}$ and $H_{n}$ are subgraphs with 16 nodes, which is known to be difficult to discriminate by standard GNNs.
The first dataset use $H_{p}$ and $H_{n}$, by which positive and negative classes are defined. We first generated a random connected graph, connected with one of $H_{p}$ or $H_{n}$ randomly. The initial random graph has node labels ‘0’, and $H_{p}$ and $H_{n}$ have node labels ‘1’. We generated $300$ instances for each of the positive and the negative classes. We call this dataset ‘Cycle’. See Appendix A.1 for further details of the initial random graph.
In the second data, the positive and the negative classes are defined by the XOR rule, i.e., a nonlinear rule, of $H_{p}$ and $H_{n}$, which we call ‘Cycle_XOR’ By using the same random graph as the Cycle dataset, $H_{p}$ and/or $H_{n}$ are embedded into input graphs. Therefore, if a graph have one of $H_{p}$ or $H_{n}$, the class label is $y=1$, otherwise $y=0$, i.e., in the case that both of $H_{p}$ and $H_{n}$ are included or neither of them are included. We generated $150$ instances each of four states of XOR, which results in $300$ instances each of the positive and the negative classes. However, in this setting, the difference in the number of the nodes labeled as ‘1’ may make the discrimination easier (32 if both of $H_{p}$ and $H_{n}$ exist, 16 if one of $H_{p}$ and $H_{n}$ exists, and 0 if neither of $H_{p}$ and $H_{n}$ exists). To avoid this, we added an simple subgraph $H_{\rm padding}$ so that the number of the nodes labeled as ‘1’ is 32 for all graphs.
Benchmark Datasets.
As benchmark datasets, we used BZR, COX2, DHFR, ENZYMES, and PTC MR from (Morris et al.,, 2020), and ToxCast from (Wu et al.,, 2018).
Compared Methods.
For performance comparison, we used GNN methods such as GCN (Kipf and Welling,, 2017), GAT (Velickovic et al.,, 2018), GATv2 (Brody et al.,, 2022), GIN (Xu et al.,, 2019), PNA (Corso et al.,, 2020), GIN-AK (Zhao et al.,, 2022), PPGN (Maron et al., 2019a, ), and ESC-GNN (Yan et al.,, 2024). For node attributes, EIN only used discrete node labels, while compared GNN methods also incorporate continuous node attributes. GCN, GAT, GATv2, and GIN optimizes the number of the units $\{64,128,256\}$ and the number of epoch by the validation set, and other settings follow (You et al.,, 2021). PNA optimizes the number of the units $\{16,32,64\}$ and the number of epochs by the validation set, and the number of the message passing is set as $2$. Other settings of GIN-AK, PPGN, and ESC-GNN follow the author implementation. For EIN, $\text{maxpat}\in\{5,10\}$, $K\in\{2,6,10\}$, and $\tau_{\rm max}\in\{1,30\}$ for $\boldsymbol{\theta}$ ($\tau_{\rm max}$ for $\boldsymbol{b}$ is fixed as $1$) are selected by the validation performance. We also evaluate performance of the combination of EIN and GNN described in § 2.4, in which GIN is employed as the combined GNN. In the results, the single EIN is denoted as EIN-$L_{0,2}$ and the EIN combined with GIN is denoted as EIN-$L_{0,2}$-GNN. For both of EIN-$L_{0,2}$ and EIN-$L_{0,2}$-GNN, the candidate of the sparsity $s$ is ${\mathcal{S}}=\{1,5,10,25,50,75,100\}$. For other details, see Appendix A.2.
4.1 Prediction Accuracy Comparison
Table 1: Accuracy comparison on synthetic and benchmark datasets (mean and std of 10 trials). The bold-face indicates that the result is comparable with the best average method in a sense of one-sided $t$-test (significance level 5%).
Method BZR COX2 DHFR ENZYMES PTC MR ToxCast Cycle Cycle_XOR
GCN 81.6 $\pm$ 2.3 78.9 $\pm$ 1.1 69.5 $\pm$ 3.3 75.0 $\pm$ 4.6 57.8 $\pm$ 6.5 59.3 $\pm$ 1.9 47.7 $\pm$ 1.7 49.9 $\pm$ 3.3
GAT 82.1 $\pm$ 3.3 77.6 $\pm$ 2.6 71.8 $\pm$ 2.0 73.2 $\pm$ 3.9 57.1 $\pm$ 5.4 60.7 $\pm$ 2.7 48.9 $\pm$ 2.0 51.8 $\pm$ 3.8
GATv2 82.3 $\pm$ 2.6 78.4 $\pm$ 2.3 69.8 $\pm$ 3.9 73.8 $\pm$ 4.0 60.2 $\pm$ 5.1 60.3 $\pm$ 2.6 50.9 $\pm$ 3.5 51.3 $\pm$ 3.3
GIN 80.9 $\pm$ 2.8 77.8 $\pm$ 2.6 69.1 $\pm$ 2.0 75.2 $\pm$ 6.1 55.5 $\pm$ 4.8 58.9 $\pm$ 1.9 50.3 $\pm$ 2.8 74.0 $\pm$ 2.0
PNA 82.9 $\pm$ 3.0 78.7 $\pm$ 0.7 66.0 $\pm$ 4.1 72.5 $\pm$ 7.9 59.9 $\pm$ 4.7 57.9 $\pm$ 2.8 50.3 $\pm$ 4.3 74.3 $\pm$ 1.9
PPGN 83.0 $\pm$ 3.6 79.8 $\pm$ 1.6 76.9 $\pm$ 4.0 65.8 $\pm$ 8.1 59.2 $\pm$ 7.8 60.2 $\pm$ 1.9 50.9 $\pm$ 2.2 78.5 $\pm$ 11.7
GIN-AK 84.0 $\pm$ 2.2 80.0 $\pm$ 3.3 75.6 $\pm$ 3.5 83.0 $\pm$ 8.6 57.7 $\pm$ 4.0 60.8 $\pm$ 1.7 52.9 $\pm$ 3.5 73.2 $\pm$ 4.3
ESC-GNN 85.4 $\pm$ 2.2 80.9 $\pm$ 2.7 80.0 $\pm$ 2.7 79.0 $\pm$ 5.6 56.4 $\pm$ 4.3 62.1 $\pm$ 2.9 54.0 $\pm$ 2.9 72.3 $\pm$ 2.9
EIN-$L_{0,2}$ 86.0 $\pm$ 2.4 80.5 $\pm$ 3.0 79.4 $\pm$ 3.6 59.8 $\pm$ 6.7 59.1 $\pm$ 4.5 60.0 $\pm$ 1.5 100.0 $\pm$ 0.0 100.0 $\pm$ 0.0
EIN-$L_{0,2}$-GNN 83.4 $\pm$ 2.9 81.1 $\pm$ 2.3 80.4 $\pm$ 2.4 73.3 $\pm$ 6.7 56.5 $\pm$ 4.4 61.6 $\pm$ 1.9 100.0 $\pm$ 0.0 98.5 $\pm$ 2.8
[…]
Table 1 shows classification accuracy on synthetic and benchmark datasets. The results on Cycle and Cycle_XOR indicate that GNN based methods cannot discriminates closed paths shown in Fig. 3 (a) and (b). On the other hand, in Cycle and Cycle_XOR datasets, EIN-$L_{0,2}$ and EIN-$L_{0,2}$-GNN achieved almost 100% classification. This indicates that EIN has a high discriminative ability about the subgraph structure, and flexibility that can capture a nonlinear relation.
For the other benchmark datasets, EIN-$L_{0,2}$ and/or EIN-$L_{0,2}$-GNN show superior to or comparable with other GNNs except for ENZYMES. We do not observe large differences between EIN-$L_{0,2}$ and EIN-$L_{0,2}$-GNN for most of datasets. EIN-$L_{0,2}$-GNN largely improves EIN-$L_{0,2}$ only for ENZYMES. For the ENZYMES dataset, the accuracy of EIN-$L_{0,2}$ is not high, which suggests achieving high discriminability only from SIF is difficult in this dataset. In such a case, combining EIN with GNN can be effective by incorporating message passing based feature.
Further, the number of the selected subgraphs in EIN, shown in the bottom of Table 1 is obviously small, from which we see that EIN can effectively identify a small number of important subgraphs for the prediction.
4.2 Examples of Post-hoc Analysis
(a) pat23
(b) pat0
(c) pat5
(d) pat9
(e) pat21
(f) SHAP values
Figure 4: Example of SHAP applied to an EIN prediction from the ToxCast dataset.
Figure 5: Decision tree on Cycle_XOR
Table 2: Mean pruning rates of EIN-$L_{0,2}$.
BZR COX2 DHFR ENZYMES PTC_MR ToxCast Cycle Cycle_XOR
# Traverse 1832 1727 978 9738 120 466 52 93
# all subgraphs 112319 76418 108486 1157661 61111 72734 31828 39883
Pruning rate 0.9837 0.9774 0.9910 0.9916 0.9980 0.9936 0.9983 0.9977
Table 3: Mean computational time for the entire Algorithm 1 (sec).
BZR COX2 DHFR ENZYMES PTC_MR ToxCast Cycle Cycle_XOR
Traverse times 50 59 338 1087 24 158 336 332
All times 1311 1248 5079 1942 1829 5323 40200 11563
We here show examples of the post-hoc analysis using the trained EIN. Figure 5 is an example of SHAP for the EIN prediction on a negative instance of the ToxCast dataset (Note that the sign of $f$ is flipped in the plot (f)). For SHAP, we used the python library https://shap.readthedocs.io/en/latest/, and see the document for detail. From Fig. 5, we can see that subgraphs (a), (b), (c), and (e) contribute to the negative class, while (d) contributes to the positive class. Figure 5 shows an example of a fitted decision tree to the trained EIN prediction for the Cycle_XOR training and test datasets. In Fig. 5, the top node classifies a graph that do not have both of $H_{p}$ and $H_{n}$ as $y=0$ (the top subgraph is included both in $H_{p}$ and $H_{n}$, but not in $H_{\rm padding}$), and the second and the third nodes consists of the XOR rule by the length 8 and 9 closed paths (the second node is the length 9 closed path and the third node is the length 8 closed path). Additional examples are shown in Appendix A.3.
4.3 Pruning Rates
Table 2 shows the mean number of the traversed nodes (mean over all $s$), and the number of all the subgraphs $|{\mathcal{H}}|$ for ${\rm maxpat}=10$. The pruning rates are defined by the ratio between them. We can obviously see that a large amount of nodes are pruned, which indicates practical effectiveness of our pruning strategy.
The computational time spent on Algorithm 1 is shown in Table 3, which is the total time for solving (2) for all $\mathcal{S}$}. The results with $\mathrm{maxpat}=10$ and $\tau_{\max}=30$ for FFN is shown, which is a slower setting (other parameters are from the best setting in a sense of the validation accuracy). In the table, ‘Traverse Times’ corresponds to the Traverse function in Algorithm 1. We see that the times required for the subgraph enumerations depend on the datasets (depends on a variety of factors such as node sizes, edge sizes, and the pruning rate), but were not dominant for most of datasets. Taking into account the fact that EIN handles all the possible subgraphs with the exact matching, we consider that EIN performs highly efficient computations. Computational time comparison with other methods are also reported in Appendix A.4.
5 Conclusions
We propose Exact subgraph Isomorphism Network (EIN). EIN combines the exact subgraph enumeration, neural networks, and the $L_{0,2}$ group sparse constraint [^1] Note that, based on a similar idea, $L_{1,2}$ based formulation is also possible, for which we show some detail and a reason we employ $L_{0,2}$ in Appendix I.. We show that predictive subgraphs can be identified efficiently by pruning unnecessarily subgraphs during the IHT without sacrificing the quality of the model. Further, the theoretical convergence of our alternating optimization is also established. We demonstrated that EIN has sufficiently high prediction accuracy compared with well-known graph neural networks despite that EIN only uses a small number of selected subgraphs. Examples of post-hoc analysis by which subgraph based decision rules can be estimated in an interpretable manner are also shown. Further improvement on the scalability is one of our important future work.
Appendix A Additional Information on Experiments
A.1 Synthetic Dataset
For both Cycle and Cycle_XOR, we first generate a random graph by the following procedure. The node size is randomly chosen from $5,6,\ldots,10$, and the number of edges are at most $20$, which are also randomly generated by choosing node pairs uniformly. From $H_{p}$, $H_{n}$, or $H_{\rm padding}$, randomly selected $n_{\rm connect}\in\{3,4,5,6\}$ nodes are connected to a randomly selected node in the initial random graph.
A.2 Other Settings of EIN
We set the number of layers of FFN as $1$. We set the terminate condition in the alternating update of Algorithm 1 as that the validation loss does not improve $5$ times. In EIN-$L_{0,2}$-GNN, the message passing in GIN was set as $3$, the number of the middle unit was $16$, and the activation function was ReLU.
A.3 Other Examples of Post-hoc Analysis
Figure 6: Decision tree on ToxCast.
Figure 7: Subgraph importance estimated by RF on Cycle_XOR dataset.
Figure 6 shows an example of a fitted decision tree to the trained EIN prediction for the ToxCast training and test datasets, which provides possible decision rule behind subgraphs and the target label. As another simple example of post-hoc analysis, Fig. 7 shows subgraph importance estimated by Random Forest (RF) (Breiman,, 2001) for Cycle_XOR. We fitted RF to a set of $(\boldsymbol{\psi}_{{\mathcal{S}}}(G_{i}),f(G_{i}))$ created by the training and test datasets and the importance is evaluated by scikit-learn feature importance of RandomForestRegressor, which is based on mean decrease of impurity by that feature. Top five important subgraphs are shown in the figures. Figure 7 indicates that the closed paths of $H_{p}$ and $H_{n}$ are identified by EIN (the top subgraph is the length 9 closed path and the third subgraph is the length 8 closed path).
A.4 Computational Times
Table 4 shows the computational times of all compared methods (For EIN-$L_{0,2}$, the values are same as Table 3). Classical GNNs (GCN, GAT, GATv2, GIN, and PNA) were relatively fast, while more recent expressive GNNs (PPGN, GIN-AK, and ESC-GNN) were much slower than them. Overall, EIN-$L_{0,2}$ was often slower than classical GNNs, while it was faster or similar to those expressive GNNs.
On the other hand, EIN-$L_{0,2}$-GNN is time-consuming in the current implementation. Our Algorithm 1 employs the gradient descent for $\boldsymbol{\theta}$ (not the ‘stochastic’ gradient descent, unlike compared GNNs). This full gradient update was time consuming for the combined GNN (backpropagation of message passing) because it was performed $\tau_{\max}(=30)$ times in each iteration of alternating update that was further performed for each $s\in{\mathcal{S}}$. Examining the possibility of computational time reduction of EIN-$L_{0,2}$-GNN (e.g., by the stochastic gradient with fewer epochs) is one of our future works.
Table 4: Mean computational time of all methods (sec).
BZR COX2 DHFR ENZYMES PTC_MR ToxCast Cycle Cycle XOR
GCN 756 806 1355 426 433 2006 770 717
GAT 851 1108 1556 494 514 2394 855 975
GATv2 1420 1806 2459 841 850 4063 1470 1568
GIN 967 1175 1784 539 541 2537 878 968
PNA 1297 1366 2118 985 939 3573 1144 1503
PPGN 41874 56586 97590 25621 12763 52513 32149 66286
GIN-AK 5539 6186 7427 4295 3680 6542 5943 6909
ESC-GNN 35988 44971 69881 22519 22309 84796 47337 59356
EIN-$L_{0,2}$ 1311 1248 5079 1942 1829 5323 40200 11563
EIN-$L_{0,2}$-GNN 26850 109417 210749 24186 29841 238816 608597 530489
Appendix B SIF by Frequency
The frequency based SIF is used in (Yoshida et al.,, 2021) in the context of a subgraph pattern based distance metric learning. They define $(H\sqsubseteq G_{i})$ as the frequency of the subgraph $H$ contained in $G_{i}$, where nodes or edges among the counted subgraphs are not allowed. Then, the feature value is defined as
$\phi_{H}(G_{i})=g(\#(H\sqsubseteq G_{i})),$
where $g$ is a monotonically non-decreasing and non-negative function such as identity function $g(x)=x$ or the indicator function $g(x)=\mathbb{I}(\#(H\sqsubseteq G_{i})>0)$ (Yoshida et al., (2021) employed $g(x)=\log(1+x)$for their evaluation). They pointed out computing the frequency without overlapping is NP-complete (Schreiber and Schwöbbermeyer,, 2005), and approximate count is also provided (see (Yoshida et al.,, 2021) for detail). Our pruning theorem (Theorem 2.1) holds for both the exact $\phi_{H}(G_{i})$ and its approximation.
Appendix C Proof of Theorem 2.1
We first transform the derivatives of the loss function with respect to $\boldsymbol{\beta}_{H}$ so that it is represented through the GML intermediate variable $\boldsymbol{h}$:
$\nabla_{\boldsymbol{\beta}_{H}}\ell^{t}$ $={\frac{\partial}{\partial\boldsymbol{\beta}_{H}}}\sum_{i=1}^{n}\mathrm{loss}(y_{i},f(G_{i});\boldsymbol{B}^{t},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})$
$=\sum_{i=1}^{n}{\frac{\partial\boldsymbol{h}^{\top}}{\partial\boldsymbol{\beta}_{H}}}{\frac{\partial\mathrm{loss}(y_{i},f(G_{i});\boldsymbol{B}^{t},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})}{\partial\boldsymbol{h}}}.$
From the definition of $\boldsymbol{h}$, we see ${\frac{\partial\boldsymbol{h}^{\top}}{\partial\boldsymbol{\beta}_{H}}}=\psi_{H}(G_{i})\boldsymbol{I}_{K}$, where $\boldsymbol{I}_{K}$ is the $K\times K$ identity matrix. As a result, we have
$\nabla_{\boldsymbol{\beta}_{H}}\ell^{t}=\sum_{i=1}^{n}{\frac{\partial\mathrm{loss}(y_{i},f(G_{i});\boldsymbol{B}^{t},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})}{\partial\boldsymbol{h}}}\psi_{H}(G_{i}).$
Since $\delta^{t}_{ik}={\frac{\partial\mathrm{loss}(y_{i},f(G_{i});\boldsymbol{B}^{t},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})}{\partial h_{k}}}$,
$\|\nabla_{\boldsymbol{\beta}_{H^{\prime}}}\ell^{t}\|_{2}=\sqrt{\sum_{k=1}^{K}\left(\sum_{i=1}^{n}\delta^{t}_{ik}\psi_{H^{\prime}}(G_{i})\right)^{2}}.$ (12)
From the definition, $0\leq\psi_{H^{\prime}}(G_{i})\leq\psi_{H}(G_{i})$. Then, the upper and lower bound of the inner sum of (12), i.e., $\sum_{i=1}^{n}\delta^{t}_{ik}\psi_{H^{\prime}}(G_{i})$, can be derived as
$\sum_{i=1}^{n}\delta^{t}_{ik}\psi_{H^{\prime}}(G_{i})\leq\sum_{\{i\mid\delta^{t}_{ik}>0\}}\delta^{t}_{ik}\psi_{H^{\prime}}(G_{i})\leq\sum_{\{i\mid\delta^{t}_{ik}>0\}}\delta^{t}_{ik}\psi_{H}(G_{i}),$
$\sum_{i=1}^{n}\delta^{t}_{ik}\psi_{H^{\prime}}(G_{i})\geq\sum_{\{i\mid\delta^{t}_{ik}<0\}}\delta^{t}_{ik}\psi_{H^{\prime}}(G_{i})\geq\sum_{\{i\mid\delta^{t}_{ik}<0\}}\delta^{t}_{ik}\psi_{H}(G_{i}).$
Therefore,
$\left(\sum_{i=1}^{n}\delta^{t}_{ik}\psi_{H^{\prime}}(G_{i})\right)^{2}$
$\leq\max\left\{\left(\sum_{\{i\mid\delta^{t}_{ik}>0\}}\delta^{t}_{ik}\psi_{H}(G_{i})\right)^{2},\ \left(\sum_{\{i\mid\delta^{t}_{ik}<0\}}\delta^{t}_{ik}\psi_{H}(G_{i})\right)^{2}\right\},$
which results in (4).
Appendix D Lower Bound with respect to Step Length
$\mathop{\text{argmin}}_{\gamma\in\Gamma}\frac{1}{\gamma}\|\boldsymbol{\beta}_{H}^{t}-\gamma\nabla_{\boldsymbol{\beta}_{H}}\ell^{t}\|_{2}$
$=\mathop{\text{argmin}}_{\gamma\in\Gamma}\frac{1}{\gamma^{2}}\|\boldsymbol{\beta}_{H}^{t}-\gamma\nabla_{\boldsymbol{\beta}_{H}}\ell^{t}\|_{2}^{2}$
$=\mathop{\text{argmin}}_{\gamma\in\Gamma}\|\boldsymbol{\beta}_{H}^{t}\|_{2}^{2}\frac{1}{\gamma^{2}}-2\left\langle\boldsymbol{\beta}_{H}^{t},\nabla_{\boldsymbol{\beta}_{H}}\ell^{t}\right\rangle\frac{1}{\gamma}$
$=\mathop{\text{argmin}}_{\gamma\in\Gamma}\|\boldsymbol{\beta}_{H}^{t}\|_{2}^{2}\left(\frac{1}{\gamma}-\frac{\left\langle\boldsymbol{\beta}_{H}^{t},\nabla_{\boldsymbol{\beta}_{H}}\ell^{t}\right\rangle}{\|\boldsymbol{\beta}_{H}^{t}\|_{2}^{2}}\right)^{2}$
Because of $\gamma>0$, if $\left\langle\boldsymbol{\beta}_{H}^{t},\nabla_{\boldsymbol{\beta}_{H}}\ell^{t}\right\rangle>0$, we see that
$\gamma=\frac{\|\boldsymbol{\beta}_{H}^{t}\|_{2}^{2}}{\left\langle\boldsymbol{\beta}_{H}^{t},\nabla_{\boldsymbol{\beta}_{H}}\ell^{t}\right\rangle}$
is the minimizer. On the other hand, if $\left\langle\boldsymbol{\beta}_{H}^{t},\nabla_{\boldsymbol{\beta}_{H}}\ell^{t}\right\rangle\leq 0$, the minimizer is $\gamma=\max\Gamma$.
Further, in practice, if $\Gamma$ is fixed beforehand, it is also possible to simply calculate $\frac{1}{\gamma}\|\boldsymbol{\beta}_{H}^{t}-\gamma\nabla_{\boldsymbol{\beta}_{H}}\ell^{t}\|_{2}$ for all $\gamma\in\Gamma$.
Appendix E Proof of Lemma 2.1
Proof.
First, we consider the update of $\boldsymbol{B}$ based on (Damadi and Shen,, 2022). For $\boldsymbol{B}$, we assume that $\gamma^{t}_{\boldsymbol{B}}$ is sufficiently small such that $L\leq\rho/\gamma^{t}_{\boldsymbol{B}}<1/\gamma^{t}_{\boldsymbol{B}}$, which can be satisfied by finite iterations of the backtrack search.
It is known that $L$-smoothness (10) is equivalent to
$\ell(\boldsymbol{B}^{\prime},\boldsymbol{b}^{\prime},\boldsymbol{\theta}^{\prime})\leq\ell(\boldsymbol{B},\boldsymbol{b},\boldsymbol{\theta})+\left\langle{\frac{\partial\ell(\boldsymbol{B},\boldsymbol{b},\boldsymbol{\theta})}{\partial\boldsymbol{z}}},\boldsymbol{z}^{\prime}-\boldsymbol{z}\right\rangle+\frac{L}{2}\|\boldsymbol{z}^{\prime}-\boldsymbol{z}\|_{2}^{2}$ (13)
for $\|\boldsymbol{B}\|_{0}\leq s$ and $\|\boldsymbol{B}^{\prime}\|_{0}\leq s$. From (13),
$\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})$ $\leq\ell(\boldsymbol{B}^{t},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})+\left\langle\nabla_{\boldsymbol{B}}\ell^{t},\boldsymbol{B}^{t+1}-\boldsymbol{B}^{t}\right\rangle+\frac{L}{2}\|\boldsymbol{B}^{t+1}-\boldsymbol{B}^{t}\|_{F}^{2}$ (14)
where $\|\boldsymbol{B}\|_{0}\leq s$ and $\|\boldsymbol{B}^{\prime}\|_{0}\leq s$. Let
$S$ $=S_{t+1}\cup S_{t},$
$\overline{S}_{t+1}$ $=S\setminus S_{t+1},$
$\overline{S}_{t}$ $=S\setminus S_{t}.$
When these indices are used as a subscript of $\boldsymbol{B}$, it represents a submatrix of $\boldsymbol{B}$ consisting of columns specified by these indices. From the definition of the hard thresholding $H_{s}$,
$\boldsymbol{B}^{t+1}_{S_{t+1}}$ $=\boldsymbol{B}^{t+1}_{S_{t+1}}-\gamma^{t}_{\boldsymbol{B}}\nabla_{\boldsymbol{B}}\ell^{t}_{S_{t+1}},$
$\boldsymbol{B}^{t+1}_{\overline{S}_{t+1}}$ $=\boldsymbol{0}.$
By combining these with (14),
$\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})-\ell(\boldsymbol{B}^{t},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})$
$\leq\left\langle\nabla_{\boldsymbol{B}}\ell^{t}_{S},\boldsymbol{B}^{t+1}_{S}-\boldsymbol{B}^{t}_{S}\right\rangle+\frac{L}{2}\|\boldsymbol{B}^{t+1}_{S}-\boldsymbol{B}^{t}_{S}\|_{F}^{2}$
$=\left\langle\nabla_{\boldsymbol{B}}\ell^{t}_{S_{t+1}},\boldsymbol{B}^{t+1}_{S_{t+1}}-\boldsymbol{B}^{t}_{S_{t+1}}\right\rangle_{F}-\left\langle\nabla_{\boldsymbol{B}}\ell^{t}_{\overline{S}_{t+1}},\boldsymbol{B}^{t}_{\overline{S}_{t+1}}\right\rangle_{F}$
$\qquad+\frac{L}{2}\|\boldsymbol{B}^{t+1}_{S_{t+1}}-\boldsymbol{B}^{t}_{S_{t+1}}\|^{2}_{F}+\frac{L}{2}\|\boldsymbol{B}^{t}_{\overline{S}_{t+1}}\|^{2}_{F}$
$=-\gamma^{t}_{\boldsymbol{B}}\|\nabla_{\boldsymbol{B}}\ell^{t}_{S_{t+1}}\|_{F}^{2}-\left\langle\nabla_{\boldsymbol{B}}\ell^{t}_{\overline{S}_{t+1}},\boldsymbol{B}^{t}_{\overline{S}_{t+1}}\right\rangle_{F}$
$\qquad+\frac{L}{2}(\gamma^{t}_{\boldsymbol{B}})^{2}\|\nabla_{\boldsymbol{B}}\ell^{t}_{S_{t+1}}\|_{F}^{2}+\frac{L}{2}\|\boldsymbol{B}^{t}_{\overline{S}_{t+1}}\|^{2}_{F}$
$=-\left(1-\frac{\gamma^{t}_{\boldsymbol{B}}L}{2}\right)\gamma^{t}_{\boldsymbol{B}}\|\nabla_{\boldsymbol{B}}\ell^{t}_{S_{t+1}}\|_{F}^{2}+\frac{L}{2}\|\boldsymbol{B}^{t}_{\overline{S}_{t+1}}\|^{2}_{F}-\left\langle\nabla_{\boldsymbol{B}}\ell^{t}_{\overline{S}_{t+1}},\boldsymbol{B}^{t}_{\overline{S}_{t+1}}\right\rangle_{F}$ (15)
By using $L\leq 1/\gamma_{\boldsymbol{B}}^{t}$, for the second and the third terms of (15),
$\frac{L}{2}\|\boldsymbol{B}^{t}_{\overline{S}_{t+1}}\|^{2}_{F}-\left\langle\nabla_{\boldsymbol{B}}\ell^{t}_{\overline{S}_{t+1}},\boldsymbol{B}^{t}_{\overline{S}_{t+1}}\right\rangle_{F}$
$\leq\frac{1}{2\gamma^{t}_{\boldsymbol{B}}}\|\boldsymbol{B}^{t}_{\overline{S}_{t+1}}\|^{2}_{F}-\left\langle\nabla_{\boldsymbol{B}}\ell^{t}_{\overline{S}_{t+1}},\boldsymbol{B}^{t}_{\overline{S}_{t+1}}\right\rangle_{F}$
$=\frac{1}{2\gamma^{t}_{\boldsymbol{B}}}\left\|\boldsymbol{B}^{t}_{\overline{S}_{t+1}}-\gamma^{t}_{\boldsymbol{B}}\nabla_{\boldsymbol{B}}\ell^{t}_{\overline{S}_{t+1}}\right\|_{F}^{2}-\frac{\gamma^{t}_{\boldsymbol{B}}}{2}\left\|\nabla_{\boldsymbol{B}}\ell^{t}_{\overline{S}_{t+1}}\right\|_{F}^{2}$ (16)
From the definition of the hard thresholding $H_{s}$, for $\forall p\in\overline{S}_{t+1}$ and $\forall q\in\overline{S}_{t}$, we see
$\|\boldsymbol{B}^{t}_{p}-\gamma^{t}_{\boldsymbol{B}}\nabla_{\boldsymbol{B}}\ell^{t}_{p}\|_{2}$ $\leq\|\boldsymbol{B}^{t}_{q}-\gamma^{t}_{\boldsymbol{B}}\nabla_{\boldsymbol{B}}\ell^{t}_{q}\|_{2}$
$=\gamma^{t}_{\boldsymbol{B}}\|\nabla_{\boldsymbol{B}}\ell^{t}_{q}\|_{2}$
Since
$\left|\overline{S}_{t+1}\right|=s-\left|\overline{S}_{t+1}\cap\overline{S}_{t}\right|=\left|\overline{S}_{t}\right|,$
we obtain an upper bound of the first term of (16) as
$\|\boldsymbol{B}^{t}_{\overline{S}_{t+1}}-\gamma^{t}_{\boldsymbol{B}}\nabla_{\boldsymbol{B}}\ell^{t}_{\overline{S}_{t+1}}\|^{2}_{F}\leq(\gamma^{t}_{\boldsymbol{B}})^{2}\|\nabla_{\boldsymbol{B}}\ell^{t}_{\overline{S}_{t}}\|^{2}_{F}.$ (17)
By applying (17) and (16) to (15),
$\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})-\ell(\boldsymbol{B}^{t},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})$
$\leq-\left(1-\frac{\gamma^{t}_{\boldsymbol{B}}L}{2}\right)\gamma^{t}_{\boldsymbol{B}}\|\nabla_{\boldsymbol{B}}\ell^{t}_{S_{t+1}}\|_{F}^{2}+\frac{\gamma^{t}_{\boldsymbol{B}}}{2}\|\nabla_{\boldsymbol{B}}\ell^{t}_{\overline{S}_{t}}\|^{2}_{F}-\frac{\gamma^{t}_{\boldsymbol{B}}}{2}\|\nabla_{\boldsymbol{B}}\ell^{t}_{\overline{S}_{t+1}}\|_{F}^{2}$
$=-\left(1-\frac{\gamma^{t}_{\boldsymbol{B}}L}{2}\right)\gamma^{t}_{\boldsymbol{B}}\|\nabla_{\boldsymbol{B}}\ell^{t}_{S_{t}\cap S_{t+1}}\|_{F}^{2}-\left(\frac{1}{2}-\frac{\gamma^{t}_{\boldsymbol{B}}L}{2}\right)\gamma^{t}_{\boldsymbol{B}}\|\nabla_{\boldsymbol{B}}\ell^{t}_{\overline{S_{t}}}\|_{F}^{2}$
$\qquad-\frac{\gamma^{t}_{\boldsymbol{B}}}{2}\|\nabla_{\boldsymbol{B}}\ell^{t}_{\overline{S}_{t+1}}\|_{F}^{2}$
$\leq-\left(\frac{1}{2}-\frac{\gamma^{t}_{\boldsymbol{B}}L}{2}\right)\gamma^{t}_{\boldsymbol{B}}\left(\|\nabla_{\boldsymbol{B}}\ell^{t}_{S_{t}\cap S_{t+1}}\|_{F}^{2}+\|\nabla_{\boldsymbol{B}}\ell^{t}_{\overline{S_{t}}}\|_{F}^{2}+\|\nabla_{\boldsymbol{B}}\ell^{t}_{\overline{S}_{t+1}}\|_{F}^{2}\right)$
$=-\left(1-\gamma^{t}_{\boldsymbol{B}}L\right)\frac{\gamma^{t}_{\boldsymbol{B}}}{2}\|\nabla_{\boldsymbol{B}}\ell^{t}_{S}\|_{F}^{2}.$
Further, using $L\leq\rho/\gamma^{t}_{\boldsymbol{B}}$, we obtain
$\ell(\boldsymbol{B}^{t},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})-\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})$ $\geq\left(1-\gamma^{t}_{\boldsymbol{B}}L\right)\frac{\gamma^{t}_{\boldsymbol{B}}}{2}\|\nabla_{\boldsymbol{B}}\ell^{t}_{S}\|_{F}^{2}$
$\geq\left(1-\rho\right)\frac{\gamma^{t}_{\boldsymbol{B}}}{2}\|\nabla_{\boldsymbol{B}}\ell^{t}_{S}\|_{F}^{2}$
$\geq c_{\boldsymbol{B}}\left(1-\rho\right)\frac{\gamma^{t}_{\boldsymbol{B}}}{2}\|\nabla_{\boldsymbol{B}}\ell^{t}_{S}\|_{F}^{2}.$ (18)
As a result, we see that (7) can be satisfied by a sufficiently small $\gamma_{\boldsymbol{B}}^{t}$. Further, since (7) holds when $L\leq\rho/\gamma_{\boldsymbol{B}}^{t}$, which means $\gamma_{\boldsymbol{B}}^{t}\leq\rho/L$. We also see $\gamma_{\boldsymbol{B}}^{t}\geq\rho^{2}/L$ (the backtrack stops before $\gamma_{\boldsymbol{B}}^{t}$ becomes smaller than $\rho^{2}/L$).
For $\boldsymbol{b}$, the condition (8) can be directly obtained from $L$-smoothness and the assumption $L\leq 1/\gamma_{\boldsymbol{b}}^{t,\tau}$ as
$\ell(\boldsymbol{B}^{t},\boldsymbol{b}^{t,\tau+1},\boldsymbol{\theta}^{t})$ $\leq\ell(\boldsymbol{B}^{t},\boldsymbol{b}^{t,\tau},\boldsymbol{\theta}^{t})+\left\langle\nabla_{\boldsymbol{b}}\ell^{t,\tau},\boldsymbol{b}^{t,\tau+1}-\boldsymbol{b}^{t,\tau}\right\rangle$
$\qquad+\frac{1}{2\gamma^{t,\tau}_{\boldsymbol{b}}}\|\boldsymbol{b}^{t,\tau+1}-\boldsymbol{b}^{t,\tau}\|^{2}_{2}$
$=\ell(\boldsymbol{B}^{t},\boldsymbol{b}^{t,\tau},\boldsymbol{\theta}^{t})-\gamma^{t,\tau}_{\boldsymbol{b}}\left\|\nabla_{\boldsymbol{b}}\ell^{t,\tau}\right\|_{2}^{2}+\frac{\gamma^{t,\tau}_{\boldsymbol{b}}}{2}\left\|\nabla_{\boldsymbol{b}}\ell^{t,\tau}\right\|_{2}^{2},$
from which we see
$\ell(\boldsymbol{B}^{t},\boldsymbol{b}^{t,\tau+1},\boldsymbol{\theta}^{t})-\ell(\boldsymbol{B}^{t},\boldsymbol{b}^{t,\tau},\boldsymbol{\theta}^{t})\geq\frac{\gamma^{t,\tau}_{\boldsymbol{b}}}{2}\left\|\nabla_{\boldsymbol{b}}\ell^{t,\tau}\right\|_{2}^{2}\geq c_{\boldsymbol{b}}\frac{\gamma^{t,\tau}_{\boldsymbol{b}}}{2}\left\|\nabla_{\boldsymbol{b}}\ell^{t,\tau}\right\|_{2}^{2}$
For $\boldsymbol{\theta}$, (9) can be obtained by the same procedure. We also see $\gamma_{\boldsymbol{b}}^{t,\tau}\geq\rho/L$ and $\gamma_{\boldsymbol{\theta}}^{t,\tau}\geq\rho/L$ because of the same reason as $\gamma_{\boldsymbol{B}}^{t}\geq\rho^{2}/L$. ∎
Appendix F Definition of Subdifferential
Let $f:\mathbb{R}^{d}\to\mathbb{R}\cup\{\infty\}$ be a proper and lower semicontinous function. Fréchet subdifferential of $f$ is
$\hat{\partial}f(\boldsymbol{x})=\left\{\boldsymbol{g}\mid\lim_{\boldsymbol{y}\to\boldsymbol{x}}\inf_{\boldsymbol{y}\neq\boldsymbol{x}}\frac{f(\boldsymbol{y})-f(\boldsymbol{x})-\left\langle\boldsymbol{g},\boldsymbol{y}-\boldsymbol{x}\right\rangle}{\|\boldsymbol{y}-\boldsymbol{x}\|_{2}}\geq 0\right\}.$
If $\boldsymbol{x}\notin\mathrm{dom}f$, then $\hat{\partial}f(\boldsymbol{x})=\emptyset$. If $f$ is a $C^{1}$ function, $\hat{\partial}f(\boldsymbol{x})$ coincides with the usual gradient. However, it is known that, under this definition, $\hat{\partial}f(\boldsymbol{x})$ becomes unnaturally empty in some cases (e.g., f(x)=|x|f(x)=-|x| at $x=0$), and also $\hat{\partial}f(\boldsymbol{x})$ is not a closed mapping. The limiting subdifferential of $f$ is defined through the limit values of Fréchet subdifferential:
$\partial f(\boldsymbol{x})=\left\{\boldsymbol{g}\mid\exists\boldsymbol{x}^{k}\to\boldsymbol{x},f(\boldsymbol{x}^{k})\to f(\boldsymbol{x}),\boldsymbol{g}^{k}\in\hat{\partial}f(\boldsymbol{x}^{k})\to\boldsymbol{g}\right\}.$
The limiting subdifferential satisfies Fermat’s rule, i.e., $\partial f(\boldsymbol{x})\ni\boldsymbol{0}$ if $\boldsymbol{x}$ is a local minimum, and is widely used for analysis of a nonsmooth nonconvex optimization.
Appendix G Proof of Theorem 2.2
G.1 Auxiliary Lemmas
To prove Theorem 2.2, we first prove the auxiliary lemmas G.1, G.2, G.3, and G.4
Lemma G.1.
There exists a constant $C_{1}$ that satisfies
$\ell(\boldsymbol{B}^{t},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})-\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t+1},\boldsymbol{\theta}^{t+1})\geq C_{1}\left\|\boldsymbol{z}^{t+1}-\boldsymbol{z}^{t}\right\|_{2}^{2}.$ (19)
Proof.
The left-hand-side of (19) can be decomposed as
$\ell(\boldsymbol{B}^{t},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})-\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t+1},\boldsymbol{\theta}^{t+1})$
$=\left(\ell(\boldsymbol{B}^{t},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})-\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})\right)$
$\qquad+\left(\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})-\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t+1},\boldsymbol{\theta}^{t})\right)$
$\qquad+\left(\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t+1},\boldsymbol{\theta}^{t})-\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t+1},\boldsymbol{\theta}^{t+1})\right).$ (20)
We derive lower bounds of each one of the three terms.
Consider decomposing the update of $\boldsymbol{B}$ into $S_{t}{\cap}S_{t+1}$, $\overline{S}_{t}$, and $\overline{S}_{t+1}$:
$\|\boldsymbol{B}^{t+1}-\boldsymbol{B}^{t}\|_{F}^{2}$ $=\|H_{s}(\boldsymbol{B}^{t}_{S}-\gamma_{\boldsymbol{B}}^{t}\nabla_{\boldsymbol{B}}L^{t}_{S})-\boldsymbol{B}_{S}^{t}\|_{F}^{2}$
$=\|\boldsymbol{B}^{t}_{S_{t}{\cap}S_{t+1}}-\gamma_{\boldsymbol{B}}^{t}\nabla_{\boldsymbol{B}}L^{t}_{S_{t}{\cap}S_{t+1}}-\boldsymbol{B}^{t}_{S_{t}{\cap}S_{t+1}}\|_{F}^{2}$
$\qquad+\|\gamma_{\boldsymbol{B}}^{t}\nabla_{\boldsymbol{B}}L^{t}_{\overline{S}_{t}}\|_{F}^{2}+\|\boldsymbol{B}^{t}_{\overline{S}_{t+1}}\|^{2}_{F}$
$=\|\gamma_{\boldsymbol{B}}^{t}\nabla_{\boldsymbol{B}}L^{t}_{S_{t}{\cap}S_{t+1}}\|_{F}^{2}+\|\gamma_{\boldsymbol{B}}^{t}\nabla_{\boldsymbol{B}}L^{t}_{\overline{S}_{t}}\|_{F}^{2}+\|\boldsymbol{B}^{t}_{\overline{S}_{t+1}}\|^{2}_{F}$
An upper bound of the last term is
$\|\boldsymbol{B}^{t}_{\overline{S}_{t+1}}\|_{F}$ $\leq\|\boldsymbol{B}^{t}_{\overline{S}_{t+1}}-\gamma_{\boldsymbol{B}}^{t}\nabla_{\boldsymbol{B}}\ell^{t}_{\overline{S}_{t+1}}\|_{F}+\|\gamma_{\boldsymbol{B}}^{t}\nabla_{\boldsymbol{B}}\ell^{t}_{\overline{S}_{t+1}}\|_{F}$
$\leq\gamma^{t}_{\boldsymbol{B}}\|\nabla_{\boldsymbol{B}}\ell^{t}_{\overline{S}_{t}}\|_{F}+\|\gamma_{\boldsymbol{B}}^{t}\nabla_{\boldsymbol{B}}\ell^{t}_{\overline{S}_{t+1}}\|_{F}.$
The last inequality is from (17). By the above inequality, we can further create an upper bound of $\|\boldsymbol{B}^{t+1}-\boldsymbol{B}^{t}\|_{F}^{2}$ as
$\|\boldsymbol{B}^{t+1}-\boldsymbol{B}^{t}\|_{F}^{2}$ $\leq\|\gamma_{\boldsymbol{B}}^{t}\nabla_{\boldsymbol{B}}L^{t}_{S_{t}{\cap}S_{t+1}}\|_{F}^{2}+\|\gamma_{\boldsymbol{B}}^{t}\nabla_{\boldsymbol{B}}L^{t}_{\overline{S}_{t}}\|_{F}^{2}$
$\qquad+\left(\gamma^{t}_{\boldsymbol{B}}\|\nabla_{\boldsymbol{B}}L^{t}_{\overline{S}_{t}}\|_{F}+\|\gamma_{\boldsymbol{B}}^{t}\nabla_{\boldsymbol{B}}L^{t}_{\overline{S}_{t+1}}\|_{F}\right)^{2}$
$\leq\|\gamma_{\boldsymbol{B}}^{t}\nabla_{\boldsymbol{B}}L^{t}_{S_{t}{\cap}S_{t+1}}\|_{F}^{2}+\|\gamma_{\boldsymbol{B}}^{t}\nabla_{\boldsymbol{B}}L^{t}_{\overline{S}_{t}}\|_{F}^{2}$
$\qquad+2(\gamma^{t}_{\boldsymbol{B}})^{2}\|\nabla_{\boldsymbol{B}}L^{t}_{\overline{S}_{t}}\|_{F}^{2}+2(\gamma_{\boldsymbol{B}}^{t})^{2}\|\nabla_{\boldsymbol{B}}L^{t}_{\overline{S}_{t+1}}\|_{F}^{2}$
$\leq(\gamma^{t}_{\boldsymbol{B}})^{2}\|\nabla_{\boldsymbol{B}}L^{t}_{S}\|_{F}^{2}+4(\gamma^{t}_{\boldsymbol{B}})^{2}\|\nabla_{\boldsymbol{B}}L^{t}_{{S}}\|_{F}^{2}$
$=5(\gamma^{t}_{\boldsymbol{B}})^{2}\|\nabla_{\boldsymbol{B}}L^{t}_{S}\|_{F}^{2}$
$\leq 5(\gamma^{t}_{\boldsymbol{B}})^{2}\frac{1}{c_{\boldsymbol{B}}(1-\rho)}\frac{2}{\gamma^{t}_{\boldsymbol{B}}}\left(\ell(\boldsymbol{B}^{t},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})-\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})\right)$ (21)
The second inequality is from the inequality of arithmetic and geometric means and the last inequality is from the backtrack condition (7).
The improvement of the objective function in the inner loop of $\boldsymbol{b}$ and $\boldsymbol{\theta}$ can be bounded by
$\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})-\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t+1},\boldsymbol{\theta}^{t})$
$=\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t,0},\boldsymbol{\theta}^{t})-\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t,\tau_{\max}},\boldsymbol{\theta}^{t})$
$=\sum_{\tau=0}^{\tau_{\max}-1}\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t,\tau},\boldsymbol{\theta}^{t})-\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t,\tau+1},\boldsymbol{\theta}^{t})$
$\geq\sum_{\tau=0}^{\tau_{\max}}c_{\boldsymbol{b}}\frac{\gamma^{t,\tau}_{\boldsymbol{b}}}{2}\left\|\nabla_{\boldsymbol{b}}\ell^{t,\tau}\right\|_{2}^{2},$ (22)
and
$\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t+1},\boldsymbol{\theta}^{t})-\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t+1},\boldsymbol{\theta}^{t+1})$
$=\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t+1},\boldsymbol{\theta}^{t,0})-\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t+1},\boldsymbol{\theta}^{t,\tau_{\max}})$
$=\sum_{\tau=0}^{\tau_{\max}-1}\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t+1},\boldsymbol{\theta}^{t,\tau})-\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t+1},\boldsymbol{\theta}^{t,\tau+1})$
$\geq\sum_{\tau=0}^{\tau_{\max}}c_{\boldsymbol{b}}\frac{\gamma^{t,\tau}_{\boldsymbol{\theta}}}{2}\left\|\nabla_{\boldsymbol{\theta}}\ell^{t,\tau}\right\|_{2}^{2}.$ (23)
Let $\gamma_{\max}=\max\Gamma$. From (22) and $\boldsymbol{b}^{t,\tau+1}=\boldsymbol{b}^{t,\tau}-\gamma^{t,\tau}_{\boldsymbol{b}}\nabla_{\boldsymbol{b}}L^{t,\tau}$,
$\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})-\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t+1},\boldsymbol{\theta}^{t})$
$\geq\sum_{\tau=0}^{\tau_{\max}-1}c_{\boldsymbol{b}}\frac{\gamma^{t,\tau}_{\boldsymbol{b}}}{2}\left\|\nabla_{\boldsymbol{b}}\ell^{t,\tau}\right\|_{2}^{2}$
$=\sum_{\tau=0}^{\tau_{\max}-1}c_{\boldsymbol{b}}\frac{1}{2\gamma^{t,\tau}_{\boldsymbol{b}}}\left\|\boldsymbol{b}^{t,\tau+1}-\boldsymbol{b}^{t,\tau}\right\|_{2}^{2}$
$\geq c_{\boldsymbol{b}}\frac{1}{2\gamma_{\max}}\sum_{\tau=0}^{\tau_{\max}-1}\left\|\boldsymbol{b}^{t,\tau+1}-\boldsymbol{b}^{t,\tau}\right\|_{2}^{2}$
$\geq c_{\boldsymbol{b}}\frac{1}{2\gamma_{\max}}\frac{1}{\tau_{\max}}\left(\sum_{\tau=0}^{\tau_{\max}-1}\left\|\boldsymbol{b}^{t,\tau+1}-\boldsymbol{b}^{t,\tau}\right\|_{2}\right)^{2}$
$\geq c_{\boldsymbol{b}}\frac{1}{2\gamma_{\max}}\frac{1}{\tau_{\max}}\left\|\sum_{\tau=0}^{\tau_{\max}-1}\left(\boldsymbol{b}^{t,\tau+1}-\boldsymbol{b}^{t,\tau}\right)\right\|_{2}^{2}$
$=c_{\boldsymbol{b}}\frac{1}{2\gamma_{\max}}\frac{1}{\tau_{\max}}\left\|\boldsymbol{b}^{t,\tau_{\max}}-\boldsymbol{b}^{t,0}\right\|_{2}^{2}$ (24)
From the fourth to the fifth line, we used a relation from the Cauchy-Schwarz inequality (for $a_{i}\in\mathbb{R}$, $\sum_{i=1}^{n}a_{i}^{2}=(\sum_{i=1}^{n}a_{i}^{2})(\sum_{i=1}^{n}1^{2})/n\geq(\sum_{i=1}^{n}(a_{i}\times 1))^{2}/n$). The same inequality for $\boldsymbol{\theta}$ is derived by using (23) and $\boldsymbol{\theta}^{t,\tau+1}=\boldsymbol{\theta}^{t,\tau}-\gamma^{t,\tau}_{\boldsymbol{b}}\nabla_{\boldsymbol{\theta}}\ell^{t,\tau}$ as
$\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t+1},\boldsymbol{\theta}^{t})-\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t+1},\boldsymbol{\theta}^{t+1})\geq c_{\boldsymbol{\theta}}\frac{1}{2\gamma_{\max}}\frac{1}{\tau_{\max}}\left\|\boldsymbol{\theta}^{t,\tau_{\max}}-\boldsymbol{\theta}^{t,0}\right\|_{2}^{2}$ (25)
By substituting (21), (24), and (25) into (20), we obtain
$\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t+1},\boldsymbol{\theta}^{t+1})-\ell(\boldsymbol{B}^{t},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})$
$\geq c_{\boldsymbol{B}}\left(1-\rho\right)\frac{1}{10\gamma^{t}_{\boldsymbol{B}}}\|\boldsymbol{B}^{t+1}-\boldsymbol{B}^{t}\|_{F}^{2}+c_{\boldsymbol{b}}\frac{1}{2\gamma_{\max}}\frac{1}{\tau_{\max}}\left\|\boldsymbol{b}^{t+1}-\boldsymbol{b}^{t}\right\|_{2}^{2}$
$\qquad+c_{\boldsymbol{\theta}}\frac{1}{2\gamma_{\max}}\frac{1}{\tau_{\max}}\left\|\boldsymbol{\theta}^{t+1}-\boldsymbol{\theta}^{t}\right\|_{2}^{2}$
$\geq C_{1}\left\|\boldsymbol{z}^{t+1}-\boldsymbol{z}^{t}\right\|_{2}^{2},$
where $C_{1}=\min\{c_{\boldsymbol{B}}\left(1-\rho\right)\frac{1}{10\gamma^{t}_{\boldsymbol{B}}},c_{\boldsymbol{b}}\frac{1}{2\gamma_{\max}}\frac{1}{\tau_{\max}},c_{\boldsymbol{\theta}}\frac{1}{2\gamma_{\max}}\frac{1}{\tau_{\max}}\}$
Lemma G.2.
For some constant $C_{\boldsymbol{b}}$ and $C_{\boldsymbol{\theta}}$,
$\left\|\nabla_{\boldsymbol{b}}\ell^{t}\right\|_{2}$ $\leq C_{\boldsymbol{b}}\|\boldsymbol{b}^{t}-\boldsymbol{b}^{t+1}\|_{2},$ (26)
$\left\|\nabla_{\boldsymbol{\theta}}\ell^{t}\right\|_{2}$ $\leq C_{\boldsymbol{\theta}}\|\boldsymbol{\theta}^{t}-\boldsymbol{\theta}^{t+1}\|_{2}.$ (27)
Proof.
From (8), we have
$\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})-\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t+1},\boldsymbol{\theta}^{t})$ $\geq\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t,0},\boldsymbol{\theta}^{t})-\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t,1},\boldsymbol{\theta}^{t})$
$\geq c_{\boldsymbol{b}}\frac{\gamma^{t,0}_{\boldsymbol{b}}}{2}\|\nabla_{\boldsymbol{b}}\ell^{t,0}\|_{2}^{2},$
and from the multi-dimensional variant of the mean value theorem, there exists $\boldsymbol{\xi}\in(\boldsymbol{b}^{t},\boldsymbol{b}^{t+1})$ that satisfies
$\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})-\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t+1},\boldsymbol{\theta}^{t})=\left\langle{\frac{\partial\ell(\boldsymbol{B}^{t+1},\boldsymbol{\xi},\boldsymbol{\theta}^{t})}{\partial\boldsymbol{b}}},\boldsymbol{b}^{t}-\boldsymbol{b}^{t+1}\right\rangle.$
Therefore, we see
$c_{\boldsymbol{b}}\frac{\gamma^{t,0}_{\boldsymbol{b}}}{2}\|\nabla_{\boldsymbol{b}}\ell^{t,0}\|_{2}^{2}$ $\leq\left\langle{\frac{\partial\ell(\boldsymbol{B}^{t+1},\boldsymbol{\xi},\boldsymbol{\theta}^{t})}{\partial\boldsymbol{b}}},\boldsymbol{b}^{t}-\boldsymbol{b}^{t+1}\right\rangle$
$\leq\left\|{\frac{\partial\ell(\boldsymbol{B}^{t+1},\boldsymbol{\xi},\boldsymbol{\theta}^{t})}{\partial\boldsymbol{b}}}\right\|_{2}d_{b},$ (28)
where $d_{b}\coloneqq\left\|\boldsymbol{b}^{t}-\boldsymbol{b}^{t+1}\right\|_{2}$. On the other hand, from $L$-smoothness, we see
$\left\|{\frac{\partial\ell(\boldsymbol{B}^{t+1},\boldsymbol{\xi},\boldsymbol{\theta}^{t})}{\partial\boldsymbol{b}}}-{\frac{\partial\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})}{\partial\boldsymbol{b}}}\right\|_{2}\leq L\left\|\boldsymbol{\xi}-\boldsymbol{b}^{t}\right\|_{2}\leq Ld_{b},$
from which we further obtain
$\left\|{\frac{\partial\ell(\boldsymbol{B}^{t+1},\boldsymbol{\xi},\boldsymbol{\theta}^{t})}{\partial\boldsymbol{b}}}\right\|_{2}$
$=\left\|{\frac{\partial\ell(\boldsymbol{B}^{t+1},\boldsymbol{\xi},\boldsymbol{\theta}^{t})}{\partial\boldsymbol{b}}}-{\frac{\partial\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})}{\partial\boldsymbol{b}}}+{\frac{\partial\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})}{\partial\boldsymbol{b}}}\right\|_{2}$
$\leq\left\|{\frac{\partial\ell(\boldsymbol{B}^{t+1},\boldsymbol{\xi},\boldsymbol{\theta}^{t})}{\partial\boldsymbol{b}}}-{\frac{\partial\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})}{\partial\boldsymbol{b}}}\right\|_{2}+\left\|{\frac{\partial\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})}{\partial\boldsymbol{b}}}\right\|_{2}$
$\leq Ld_{b}+\left\|{\frac{\partial\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})}{\partial\boldsymbol{b}}}\right\|_{2}.$ (29)
By combining (28) and (29),
$c_{\boldsymbol{b}}\frac{\gamma^{t,0}_{\boldsymbol{b}}}{2}\|\nabla_{\boldsymbol{b}}\ell^{t,0}\|_{2}^{2}$ $\leq\left(Ld_{b}+\left\|{\frac{\partial\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})}{\partial\boldsymbol{b}}}\right\|_{2}\right)d_{b}$
$=\left(Ld_{b}+\left\|\nabla_{\boldsymbol{b}}\ell^{t,0}\right\|_{2}\right)d_{b}$
from which we obtain
$Ld_{b}^{2}+\left\|\nabla_{\boldsymbol{b}}\ell^{t,0}\right\|_{2}d_{b}-c_{\boldsymbol{b}}\frac{\gamma^{t,\tau_{\max}-1}_{\boldsymbol{b}}}{2}\|\nabla_{\boldsymbol{b}}\ell^{t,0}\|_{2}^{2}\geq 0.$
By regarding the left-hand-side as a quadratic function of $d_{b}$,
$d_{b}$ $\geq\frac{-\left\|\nabla_{\boldsymbol{b}}\ell^{t,0}\right\|_{2}+\sqrt{\left\|\nabla_{\boldsymbol{b}}\ell^{t,0}\right\|_{2}^{2}+4Lc_{\boldsymbol{b}}\frac{\gamma^{t,0}_{\boldsymbol{b}}}{2}\|\nabla_{\boldsymbol{b}}\ell^{t,0}\|_{2}^{2}}}{2L}$
$=\left(\sqrt{1+2Lc_{\boldsymbol{b}}\gamma^{t,0}_{\boldsymbol{b}}}-1\right)\frac{\left\|\nabla_{\boldsymbol{b}}\ell^{t,0}\right\|_{2}}{2L}$
Therefore,
$\left\|\nabla_{\boldsymbol{b}}\ell^{t,0}\right\|_{2}$ $\leq\frac{2L}{\sqrt{1+2Lc_{\boldsymbol{b}}\gamma^{t,0}_{\boldsymbol{b}}}-1}\|\boldsymbol{b}^{t}-\boldsymbol{b}^{t+1}\|_{2}$
$\leq\frac{2L}{\sqrt{1+2\rho c_{\boldsymbol{b}}}-1}\|\boldsymbol{b}^{t}-\boldsymbol{b}^{t+1}\|_{2}.$
The last inequality is from $\gamma_{\boldsymbol{b}}^{t,0}\geq\rho/L$. The same derivation can be applied to $\boldsymbol{\theta}$. ∎
Lemma G.3.
There exists $\nabla_{\boldsymbol{z}}\tilde{\ell}(\boldsymbol{z}^{t+1})\in\partial\tilde{\ell}(\boldsymbol{z}^{t+1})$ such that
$\|\nabla_{\boldsymbol{z}}\tilde{\ell}(\boldsymbol{z}^{t+1})\|_{2}^{2}\leq C_{2}\|\boldsymbol{z}^{t+1}-\boldsymbol{z}^{t}\|_{2}^{2},$ (30)
where $C_{2}$ is a constant.
Proof.
The hard thresholding can be written as an optimization problem:
$\boldsymbol{B}^{t+1}$ $=H_{s}(\boldsymbol{B}^{t}-\gamma_{\boldsymbol{B}}\nabla_{\boldsymbol{B}}\ell^{t})$
$=\mathop{\text{argmin}}_{\boldsymbol{B}}\frac{1}{2\gamma_{\boldsymbol{B}}^{t}}\|\boldsymbol{B}-(\boldsymbol{B}^{t}-\gamma_{\boldsymbol{B}}^{t}\nabla_{\boldsymbol{B}}\ell^{t})\|_{F}^{2}+\delta_{s}(\boldsymbol{B}).$ (31)
Taking the subdifferential of (31)
$\boldsymbol{0}\in$ $\frac{1}{\gamma_{\boldsymbol{B}}^{t}}(\boldsymbol{B}^{t+1}-(\boldsymbol{B}^{t}-\gamma_{\boldsymbol{B}}^{t}\nabla_{\boldsymbol{B}}\ell^{t}))+\partial\delta_{s}(\boldsymbol{B}^{t+1})$
$-\frac{1}{\gamma_{\boldsymbol{B}}^{t}}(\boldsymbol{B}^{t+1}-(\boldsymbol{B}^{t}-\gamma_{\boldsymbol{B}}^{t}\nabla_{\boldsymbol{B}}\ell^{t}))$ $\in\partial\delta_{s}(\boldsymbol{B}^{t+1}),$
where $\partial\delta_{s}(\boldsymbol{B})$ is the subdifferential of $\delta_{s}(\boldsymbol{B})$. By using $\partial\delta_{s}(\boldsymbol{B}^{t+1})$, the subdifferential $\partial\tilde{\ell}(\boldsymbol{z}^{t+1})$ can be written as
$\partial\tilde{\ell}(\boldsymbol{z}^{t+1})$ $=\nabla_{\boldsymbol{B}}\ell^{t+1}+\partial\delta_{s}(\boldsymbol{B}^{t+1})$
$\ni\nabla_{\boldsymbol{B}}\ell^{t+1}-\frac{1}{\gamma_{\boldsymbol{B}}^{t}}(\boldsymbol{B}^{t+1}-(\boldsymbol{B}^{t}-\gamma_{\boldsymbol{B}}^{t}\nabla_{\boldsymbol{B}}\ell^{t}))$
$=\nabla_{\boldsymbol{B}}\ell^{t+1}-\nabla_{\boldsymbol{B}}\ell^{t}-\frac{1}{\gamma_{\boldsymbol{B}}^{t}}(\boldsymbol{B}^{t+1}-\boldsymbol{B}^{t})\eqqcolon\nabla_{\boldsymbol{B}}\tilde{\ell}^{t+1}$
From the triangle inequality and $L$-smoothenss,
$\|\nabla_{\boldsymbol{B}}\tilde{\ell}^{t+1}\|_{F}^{2}$ $\leq\left(\|\nabla_{\boldsymbol{B}}\ell^{t+1}-\nabla_{\boldsymbol{B}}\ell^{t}\|_{F}+\frac{1}{\gamma_{\boldsymbol{B}}^{t}}\|\boldsymbol{B}^{t+1}-\boldsymbol{B}^{t}\|_{F}\right)^{2}$
$\leq\left(L\|\boldsymbol{z}^{t+1}-\boldsymbol{z}\|_{F}+\frac{1}{\gamma_{\boldsymbol{B}}^{t}}\|\boldsymbol{B}^{t+1}-\boldsymbol{B}^{t}\|_{F}\right)^{2}$
$\leq 2L^{2}\left(\|\boldsymbol{B}^{t+1}-\boldsymbol{B}^{t}\|^{2}_{F}+\|\boldsymbol{b}^{t+1}-\boldsymbol{b}\|_{2}^{2}+\|\boldsymbol{\theta}^{t+1}-\boldsymbol{\theta}^{t}\|_{2}^{2}\right)$
$\qquad+\frac{2}{(\gamma_{\boldsymbol{B}}^{t})^{2}}\|\boldsymbol{B}^{t+1}-\boldsymbol{B}^{t}\|_{F}^{2}$
$\leq 2\left(L^{2}+\frac{1}{(\gamma_{\boldsymbol{B}}^{t})^{2}}\right)\|\boldsymbol{B}^{t+1}-\boldsymbol{B}^{t}\|_{F}^{2}$
$\qquad+2L^{2}\|\boldsymbol{b}^{t+1}-\boldsymbol{b}\|_{2}^{2}+2L^{2}\|\boldsymbol{\theta}^{t+1}-\boldsymbol{\theta}^{t}\|_{2}^{2}$ (32)
An upper bound of the derivative of $\boldsymbol{b}$ is
$\|\nabla_{\boldsymbol{b}}\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t+1},\boldsymbol{\theta}^{t+1})\|_{2}^{2}$
$\leq\bigl(\|\nabla_{\boldsymbol{b}}\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t+1},\boldsymbol{\theta}^{t+1})-\nabla_{\boldsymbol{b}}\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})\|_{2}$
$\qquad+\|\nabla_{\boldsymbol{b}}\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})\|_{2}\bigr)^{2}$
$\leq 2{L}^{2}\left(\|\boldsymbol{b}^{t+1}-\boldsymbol{b}^{t}\|_{2}^{2}+\|\boldsymbol{\theta}^{t+1}-\boldsymbol{\theta}^{t}\|_{2}^{2}\right)+2\|\nabla_{\boldsymbol{b}}\ell^{t}\|_{2}^{2}$
$\leq 2\left(L^{2}+C_{\boldsymbol{b}}^{2}\right)\|\boldsymbol{b}^{t+1}-\boldsymbol{b}^{t}\|_{2}^{2}+2{L}^{2}\|\boldsymbol{\theta}^{t+1}-\boldsymbol{\theta}^{t}\|_{2}^{2}$ (33)
The last inequality is from (26).
Similarly, for $\boldsymbol{\theta}$,
$\|\nabla_{\boldsymbol{\theta}}\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t+1},\boldsymbol{\theta}^{t+1})\|_{2}^{2}$
$\leq\bigl\{\|\nabla_{\boldsymbol{\theta}}\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t+1},\boldsymbol{\theta}^{t+1})-\nabla_{\boldsymbol{\theta}}\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t+1},\boldsymbol{\theta}^{t})\|_{2}$
$\qquad+\|\nabla_{\boldsymbol{\theta}}\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t+1},\boldsymbol{\theta}^{t})\|_{2}\bigr\}^{2}$
$\leq 2L^{2}\|\boldsymbol{\theta}^{t+1}-\boldsymbol{\theta}^{t}\|_{2}^{2}+2\|\nabla_{\boldsymbol{\theta}}\ell^{t}\|_{2}^{2}$
$\leq 2\left(L^{2}+C_{\boldsymbol{\theta}}^{2}\right)\|\boldsymbol{\theta}^{t+1}-\boldsymbol{\theta}^{t}\|_{2}^{2}$ (34)
The last inequality is from (27).
By combining (32), (33), and (34), we see
$\|\nabla_{\boldsymbol{z}}\tilde{\ell}(\boldsymbol{z}^{t+1})\|_{2}^{2}$
$=\|\nabla_{\boldsymbol{B}}\tilde{\ell}^{t+1}\|_{F}^{2}+\|\nabla_{\boldsymbol{b}}\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t+1},\boldsymbol{\theta}^{t+1})\|_{2}^{2}+\|\nabla_{\boldsymbol{\theta}}\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t+1},\boldsymbol{\theta}^{t+1})\|_{2}^{2}$
$\leq 2\left(L^{2}+\frac{1}{(\gamma_{\boldsymbol{B}}^{t})^{2}}\right)\|\boldsymbol{B}^{t+1}-\boldsymbol{B}^{t}\|_{F}^{2}+2{L}^{2}\|\boldsymbol{b}^{t+1}-\boldsymbol{b}^{t}\|_{2}^{2}+2{L}^{2}\|\boldsymbol{\theta}^{t+1}-\boldsymbol{\theta}^{t}\|_{2}^{2}$
$\qquad+2\left(L^{2}+C_{\boldsymbol{b}}^{2}\right)\|\boldsymbol{b}^{t+1}-\boldsymbol{b}^{t}\|_{2}^{2}+2L^{2}\|\boldsymbol{\theta}^{t+1}-\boldsymbol{\theta}^{t}\|_{2}^{2}$
$\qquad+2\left(L^{2}+C_{\boldsymbol{\theta}}^{2}\right)\|\boldsymbol{\theta}^{t+1}-\boldsymbol{\theta}^{t}\|_{2}^{2}$
$=2\left(L^{2}+\frac{1}{(\gamma_{\boldsymbol{B}}^{t})^{2}}\right)\|\boldsymbol{B}^{t+1}-\boldsymbol{B}^{t}\|_{F}^{2}$
$\qquad+2\left(2L^{2}+C_{\boldsymbol{b}}^{2}\right)\|\boldsymbol{b}^{t+1}-\boldsymbol{b}^{t}\|_{2}^{2}$
$\qquad+2\left(3L^{2}+C_{\boldsymbol{\theta}}^{2}\right)\|\boldsymbol{\theta}^{t+1}-\boldsymbol{\theta}^{t}\|_{2}^{2}$
$\leq C_{2}\|\boldsymbol{z}^{t+1}-\boldsymbol{z}^{t}\|_{2}^{2},$
where $C_{2}=\max\left\{2\left(L^{2}+\frac{1}{(\gamma_{\boldsymbol{B}}^{t})^{2}}\right),2\left(2L^{2}+C_{\boldsymbol{b}}^{2}\right),2\left(3L^{2}+C_{\boldsymbol{\theta}}^{2}\right)\right\}$.
Lemma G.4.
$\boldsymbol{z}^{t}$ has at least one accumulation point $z^{*}$ such that $\|\nabla_{\boldsymbol{z}}\tilde{\ell}(\boldsymbol{z}^{*})\|_{2}^{2}=0$.
Proof.
Summing (19) from $t=1$ to T1T-1, we obtain
$\ell(\boldsymbol{B}^{0},\boldsymbol{b}^{0},\boldsymbol{\theta}^{0})-\ell(\boldsymbol{B}^{T},\boldsymbol{b}^{T},\boldsymbol{\theta}^{T})\geq C_{1}\sum_{t=1}^{T}\left\|\boldsymbol{z}^{t}-\boldsymbol{z}^{t-1}\right\|_{2}^{2}$
Considering $T\to\infty$ and combining with (30)
$\infty$ $>\lim_{T\to\infty}\ell(\boldsymbol{B}^{0},\boldsymbol{b}^{0},\boldsymbol{\theta}^{0})-\ell(\boldsymbol{B}^{T},\boldsymbol{b}^{T},\boldsymbol{\theta}^{T})$
$\geq C_{1}\lim_{T\to\infty}\sum_{t=1}^{T}\left\|\boldsymbol{z}^{t}-\boldsymbol{z}^{t-1}\right\|_{2}^{2}$
$\geq\frac{C_{1}}{C_{2}}\lim_{T\to\infty}\sum_{t=1}^{T}\|\nabla_{\boldsymbol{z}}\tilde{\ell}(\boldsymbol{z}^{t})\|_{2}^{2}$
From the monotone convergence theorem, $\|\nabla_{\boldsymbol{z}}\tilde{\ell}(\boldsymbol{z}^{t})\|_{2}^{2}\to 0$. Since the sequence $z^{t}$ is assumed be bounded, we obtain the claim of the lemma. ∎
Proof of Theorem 2.2
In the proof, we use the well-known Kurdyka-Łojasiewicz property.
Definition G.1.
(Kurdyka-Łojasiewicz property) A function $\varphi(\boldsymbol{u})$ satisfies the Kurdyka-Łojasiewicz (KL) property at a point $\bar{\boldsymbol{u}}\in\mathrm{dom}(\partial\varphi)$ if there exist $\eta>0$, a neighborhood ${\mathcal{B}}_{\rho}(\bar{\boldsymbol{u}})\coloneqq\{\boldsymbol{u}\mid\|\boldsymbol{u}-\bar{\boldsymbol{u}}\|_{2}<\rho\}$, and a concave function $\phi(s)=cs^{1-\theta}$ for some $c>0$ and $\theta\in[0,1)$ such that for any $\boldsymbol{x}\in{\mathcal{B}}_{\rho}(\bar{\boldsymbol{u}})\cap\mathrm{dom}(\partial\varphi)$ and $\varphi(\bar{\boldsymbol{u}})<\varphi(\boldsymbol{u})<\varphi(\bar{\boldsymbol{u}})+\eta$, it holds
$\phi^{\prime}(|\varphi(\boldsymbol{u})-\varphi(\bar{\boldsymbol{u}})|)\mathrm{dist}(\boldsymbol{0},\partial\varphi(\boldsymbol{x}))\geq 1$
where $\mathrm{dom}(\partial\varphi)=\{\boldsymbol{u}\mid\partial\varphi(\boldsymbol{u})\neq\emptyset\}$ and $\mathrm{dist}(\boldsymbol{0},\partial\varphi(\boldsymbol{u}))=\min\{\|\boldsymbol{v}\|_{2}\mid\boldsymbol{v}\in\partial\varphi(\boldsymbol{u})\}$.
The proof of Theorem 2.2 is as follows.
Proof.
It is known that a lower semicontinuous ‘definable’ function satisfies the KL property (Bolte et al.,, 2007) (see Theorem 14 and Remark 8). Our $\ell$ is the cross-entropy loss function for a neural network with the sigmoid activation function. These components are known as definable functions (Ding and Toh,, 2025; Kranz et al.,, 2026), and the compositions of definable functions are also definable (van den Dries and Miller,, 1996; Attouch et al.,, 2010). L0 penalty $\delta_{s}$ is a semialgebratic function (Bao et al.,, 2014). A semialgebratic function is definable (Dries,, 1998), and the sum of definable functions is also definable (Attouch et al.,, 2010). As a result, $\tilde{\ell}$ is a lower semicontinuous definable function.
Here, we temporarily assume $\boldsymbol{z}^{t}\in{\mathcal{B}}_{\rho}(\boldsymbol{z}^{*})$ for $\forall t$ (we will prove the existence of such a sequence later in the proof). From the KL property,
$\phi^{\prime}(\tilde{\ell}(\boldsymbol{z})-\ell^{*})\cdot\mathrm{dist}(0,\partial\tilde{\ell}(\boldsymbol{z}))$ $\geq 1$
$\phi^{\prime}(\tilde{\ell}(\boldsymbol{z})-\ell^{*})$ $\geq\frac{1}{\mathrm{dist}(0,\partial\tilde{\ell}(\boldsymbol{z}))}\geq\frac{1}{\|{\nabla_{\boldsymbol{z}}\tilde{\ell}(\boldsymbol{z})}\|_{2}},$
where $\ell^{*}=\tilde{\ell}(\boldsymbol{z}^{*})$, $\forall\nabla_{\boldsymbol{z}}\tilde{\ell}(\boldsymbol{z})\in\partial\tilde{\ell}(\boldsymbol{z})$, and $\phi(s)=cs^{1-\theta}$ is a concave function for some $c>0$ and $\theta\in[0,1)$. Since $\phi$ is concave, $\phi(u)-\phi(v)\geq\phi^{\prime}(u)(u-v)$. Substituting $u=\tilde{\ell}(\boldsymbol{z}^{t})$-$\ell^{*}$ and $v=\tilde{\ell}(\boldsymbol{z}^{t+1})$-$\ell^{*}$ into this inequality, we obtain
$\phi(\tilde{\ell}(\boldsymbol{z}^{t})-\ell^{*})-\phi(\tilde{\ell}(\boldsymbol{z}^{t+1})-\ell^{*})$ $\geq\phi^{\prime}(\tilde{\ell}(\boldsymbol{z}^{t})-\ell^{*})\cdot(\tilde{\ell}(\boldsymbol{z}^{t})-\tilde{\ell}(\boldsymbol{z}^{t+1}))$
$\geq\frac{\tilde{\ell}(\boldsymbol{z}^{t})-\tilde{\ell}(\boldsymbol{z}^{t+1})}{\|{\nabla_{\boldsymbol{z}}\tilde{\ell}(\boldsymbol{z}^{t})}\|_{2}}.$
By using (19) of lemma G.1 and (30) of lemma G.3, there exists ${\nabla_{\boldsymbol{z}}\tilde{\ell}(\boldsymbol{z}^{t})}\in\partial\tilde{\ell}(\boldsymbol{z}^{t})$ that satisfies
$\phi(\tilde{\ell}(\boldsymbol{z}^{t})-\ell^{*})-\phi(\tilde{\ell}(\boldsymbol{z}^{t+1})-\ell^{*})\geq\frac{C_{1}\left\|\boldsymbol{z}^{t+1}-\boldsymbol{z}^{t}\right\|_{2}^{2}}{\|{\nabla_{\boldsymbol{z}}\tilde{\ell}(\boldsymbol{z}^{t})}\|_{2}}\geq\frac{C_{1}\left\|\boldsymbol{z}^{t+1}-\boldsymbol{z}^{t}\right\|_{2}^{2}}{\sqrt{C_{2}}\|\boldsymbol{z}^{t}-\boldsymbol{z}^{t-1}\|_{2}}$
By arranging this inequality,
$\left\|\boldsymbol{z}^{t+1}-\boldsymbol{z}^{t}\right\|_{2}^{2}$ $\leq\frac{\sqrt{C_{2}}}{C_{1}}\|\boldsymbol{z}^{t}-\boldsymbol{z}^{t-1}\|_{2}\left(\phi(\tilde{\ell}(\boldsymbol{z}^{t})-\ell^{*})-\phi(\tilde{\ell}(\boldsymbol{z}^{t+1})-\ell^{*})\right)$
$2\left\|\boldsymbol{z}^{t+1}-\boldsymbol{z}^{t}\right\|_{2}$ $\leq\|\boldsymbol{z}^{t}-\boldsymbol{z}^{t-1}\|_{2}+\frac{\sqrt{C_{2}}}{C_{1}}\left(\phi(\tilde{\ell}(\boldsymbol{z}^{t})-\ell^{*})-\phi(\tilde{\ell}(\boldsymbol{z}^{t+1})-\ell^{*})\right).$ (35)
To derive the last inequality, the inequality of arithmetic and geometric means is used. The sum from $t=1$ to $t=T$ becomes
$2\sum_{t=1}^{T}\left\|\boldsymbol{z}^{t+1}-\boldsymbol{z}^{t}\right\|_{2}$ $\leq\sum_{t=1}^{T}\|\boldsymbol{z}^{t}-\boldsymbol{z}^{t-1}\|_{2}$
$\qquad+\sum_{t=1}^{T}\frac{\sqrt{C_{2}}}{C_{1}}\left(\phi(\tilde{\ell}(\boldsymbol{z}^{t})-\ell^{*})-\phi(\tilde{\ell}(\boldsymbol{z}^{t+1})-\ell^{*})\right)$
$\sum_{t=1}^{T}\left\|\boldsymbol{z}^{t+1}-\boldsymbol{z}^{t}\right\|_{2}$ $\leq\|\boldsymbol{z}^{1}-\boldsymbol{z}^{0}\|_{2}$
$\qquad+\frac{\sqrt{C_{2}}}{C_{1}}\left(\phi(\tilde{\ell}(\boldsymbol{z}^{1})-\ell^{*})-\phi(\tilde{\ell}(\boldsymbol{z}^{T+1})-\ell^{*})\right)$
$\leq\|\boldsymbol{z}^{1}-\boldsymbol{z}^{0}\|_{2}+\frac{\sqrt{C_{2}}}{C_{1}}\phi(\tilde{\ell}(\boldsymbol{z}^{1})-\ell^{*})$ (36)
We see that the last upper bound does not depend on $T$, by which we obtain
$\lim_{T\to\infty}\sum_{t=1}^{T}\left\|\boldsymbol{z}^{t+1}-\boldsymbol{z}^{t}\right\|_{2}<\infty.$
Therefore, we obtain $\lim_{t\to\infty}\left\|\boldsymbol{z}^{t+1}-\boldsymbol{z}^{t}\right\|_{2}=0$, which is a necessary condition for the convergence of an infinite series. This means that $\boldsymbol{z}^{t}$ is a Cauchy sequence. From (30) of lemma G.3,
$\infty>\lim_{T\to\infty}\sum_{t=1}^{T}\left\|\boldsymbol{z}^{t+1}-\boldsymbol{z}^{t}\right\|_{2}\geq\lim_{T\to\infty}\sum_{t=1}^{T}\frac{1}{C_{2}}\|{\nabla_{\boldsymbol{z}}\tilde{\ell}(\boldsymbol{z}^{t+1})}\|_{2}^{2}.$
Here again, to assure the convergence of an infinite series to a finite value, we have
$\lim_{t\to\infty}\|\nabla_{\boldsymbol{z}}\tilde{\ell}(\boldsymbol{z}^{t})\|_{2}^{2}\to 0.$
Next, we prove our temporarily assumption $\boldsymbol{z}^{t}\in{\mathcal{B}}_{\rho}(\boldsymbol{z}^{*})$ by using similar approach to Theorem 8 of (Attouch et al.,, 2010). We re-define the index $t$ such that it satisfies the following inequality as $t=0$:
$\|\boldsymbol{z}^{*}-\boldsymbol{z}^{t}\|_{2}+2\sqrt{\frac{\tilde{\ell}(\boldsymbol{z}^{t})-\ell^{*}}{C_{1}}}+\frac{\sqrt{C_{2}}}{C_{1}}\phi(\tilde{\ell}(\boldsymbol{z}^{t})-\ell^{*})\leq\rho.$ (37)
Because of lemma G.4, there exist $t$ that satisfies the above inequality. After replacing the above $t$ with 0, obviously, $\boldsymbol{z}^{0}\in{\mathcal{B}}_{\rho}(\boldsymbol{z}^{*})$. Then, from the triangle inequality and (19) of lemma G.1, we see
$\|\boldsymbol{z}^{*}-\boldsymbol{z}^{1}\|_{2}$ $\leq\|\boldsymbol{z}^{*}-\boldsymbol{z}^{0}\|_{2}+\|\boldsymbol{z}^{0}-\boldsymbol{z}^{1}\|_{2}$
$\leq\|\boldsymbol{z}^{*}-\boldsymbol{z}^{0}\|_{2}+\sqrt{\frac{\ell(\boldsymbol{z}^{0})-\ell(\boldsymbol{z}^{1})}{C_{1}}}$
$\leq\|\boldsymbol{z}^{*}-\boldsymbol{z}^{0}\|_{2}+\sqrt{\frac{\tilde{\ell}(\boldsymbol{z}^{0})-\ell^{*}}{C_{1}}}\leq\rho.$ (38)
Therefore, by comparing the last inequality with (37), we see $\boldsymbol{z}^{1}\in{\mathcal{B}}_{\rho}(\boldsymbol{z}^{*})$. Next, we assume $\boldsymbol{z}^{k}\in{\mathcal{B}}_{\rho}(\boldsymbol{z}^{*})$ holds until some $k>1$. From the triangle inequality and (36), which can be used because now we assume $\boldsymbol{z}^{t}\in{\mathcal{B}}_{\rho}(\boldsymbol{z}^{*})$ for $t\leq k$, we obtain
$\|\boldsymbol{z}^{*}-\boldsymbol{z}^{k+1}\|_{2}$ $\leq\|\boldsymbol{z}^{k+1}-\boldsymbol{z}^{1}\|_{2}+\|\boldsymbol{z}^{*}-\boldsymbol{z}^{1}\|_{2}$
$\leq\sum_{t=1}^{k}\|\boldsymbol{z}^{t+1}-\boldsymbol{z}^{t}\|_{2}+\|\boldsymbol{z}^{*}-\boldsymbol{z}^{1}\|_{2}$
$\leq\|\boldsymbol{z}^{1}-\boldsymbol{z}^{0}\|_{2}+\frac{\sqrt{C_{2}}}{C_{1}}\phi(\tilde{\ell}(\boldsymbol{z}^{1})-\ell^{*})+\|\boldsymbol{z}^{*}-\boldsymbol{z}^{1}\|_{2}$
$\leq\sqrt{\frac{\ell(\boldsymbol{z}^{0})-\ell(\boldsymbol{z}^{1})}{C_{1}}}+\frac{\sqrt{C_{2}}}{C_{1}}\phi(\tilde{\ell}(\boldsymbol{z}^{1})-\ell^{*})$
$\qquad+\|\boldsymbol{z}^{*}-\boldsymbol{z}^{0}\|_{2}+\sqrt{\frac{\tilde{\ell}(\boldsymbol{z}^{0})-\ell^{*}}{C_{1}}}$
$\leq\|\boldsymbol{z}^{*}-\boldsymbol{z}^{0}\|_{2}+2\sqrt{\frac{\tilde{\ell}(\boldsymbol{z}^{0})-\ell^{*}}{C_{1}}}+\frac{\sqrt{C_{2}}}{C_{1}}\phi(\tilde{\ell}(\boldsymbol{z}^{0})-\ell^{*}).$
The second last inequality is from (19) of lemma G.1 and (38). Therefore, from (37), we see $\boldsymbol{z}^{k+1}\in{\mathcal{B}}_{\rho}(\boldsymbol{z}^{*})$. By induction, we obtain $\boldsymbol{z}^{t}\in{\mathcal{B}}_{\rho}(\boldsymbol{z}^{*})$ for any $t\geq 0$.
Appendix H Proof of theorem 2.3
Let $\Delta_{k}\coloneqq\sum_{t=k}^{\infty}\|\boldsymbol{z}^{t+1}-\boldsymbol{z}^{t}\|_{2}$. Note that, from the triangle inequality, $\Delta_{k}\geq\|\boldsymbol{z}^{k}-\boldsymbol{z}^{*}\|_{2}$ and $\Delta_{k-1}-\Delta_{k}=\|\boldsymbol{z}^{k}-\boldsymbol{z}^{k-1}\|_{2}$. Based on the same derivation as (36), we have
$\sum_{t=k}^{T}\left\|\boldsymbol{z}^{t+1}-\boldsymbol{z}^{t}\right\|_{2}\leq\|\boldsymbol{z}^{k}-\boldsymbol{z}^{k-1}\|_{2}+\frac{\sqrt{C_{2}}}{C_{1}}\phi(\tilde{\ell}(\boldsymbol{z}^{k})-\ell^{*}).$
Since the right-hand-side of the above inequality does not depend on $T$,
$\Delta_{k}=\lim_{T\to\infty}\sum_{t=k}^{T}\left\|\boldsymbol{z}^{t+1}-\boldsymbol{z}^{t}\right\|_{2}$ $\leq\|\boldsymbol{z}^{k}-\boldsymbol{z}^{k-1}\|_{2}+\frac{\sqrt{C_{2}}}{C_{1}}\phi(\tilde{\ell}(\boldsymbol{z}^{k})-\ell^{*})$
$=\Delta_{k-1}-\Delta_{k}+\frac{c\sqrt{C_{2}}}{C_{1}}(\tilde{\ell}(\boldsymbol{z}^{k})-\ell^{*})^{1-\theta}$ (39)
Hereafter, we set $\ell^{*}=0$ without loss of generality (Attouch and Bolte,, 2009). Then, from the KL inequality,
$\phi^{\prime}(\tilde{\ell}(\boldsymbol{z}))$ $\geq\frac{1}{\|\nabla_{\boldsymbol{z}}\tilde{\ell}(\boldsymbol{z})\|_{2}}$
$c(1-\theta)(\tilde{\ell}(\boldsymbol{z}))^{-\theta}$ $\geq\frac{1}{\|\nabla_{\boldsymbol{z}}\tilde{\ell}(\boldsymbol{z})\|_{2}}$
$(\tilde{\ell}(\boldsymbol{z}))^{\theta}$ $\leq c(1-\theta)\|\nabla_{\boldsymbol{z}}\tilde{\ell}(\boldsymbol{z})\|_{2}$ (40)
Substituting this into (39) and combining with (30) of lemma G.3, we see
$\Delta_{k}$ $\leq\Delta_{k-1}-\Delta_{k}+\frac{c\sqrt{C_{2}}}{C_{1}}(c(1-\theta)\|\nabla_{\boldsymbol{z}}\tilde{\ell}(\boldsymbol{z}^{k})\|_{2})^{\frac{1-\theta}{\theta}}$
$\leq\Delta_{k-1}-\Delta_{k}+\frac{c\sqrt{C_{2}}}{C_{1}}(cC_{2}(1-\theta))^{\frac{1-\theta}{\theta}}\|\boldsymbol{z}^{k}-\boldsymbol{z}^{k-1}\|_{2}^{\frac{1-\theta}{\theta}}$
$=\Delta_{k-1}-\Delta_{k}+\frac{c\sqrt{C_{2}}}{C_{1}}(cC_{2}(1-\theta))^{\frac{1-\theta}{\theta}}(\Delta_{k-1}-\Delta_{k})^{\frac{1-\theta}{\theta}}.$ (41)
When $\theta\in(1/2,1)$, we have $\frac{1-\theta}{\theta}<1$. Further, when $k\to\infty$, we have $\Delta_{k-1}-\Delta_{k}\to+0$. Then, there exists $N_{1}$ such that $\Delta_{k-1}-\Delta_{k}\leq(\Delta_{k-1}-\Delta_{k})^{\frac{1-\theta}{\theta}}$ holds for $\forall k\geq N_{1}$. Therefore, there exists a constant $Z_{1}$ that satisfies
$\Delta_{k}^{\frac{\theta}{1-\theta}}\leq Z_{1}(\Delta_{k-1}-\Delta_{k})$
for $k\geq N_{1}$. Define $h(s)=s^{-\frac{\theta}{1-\theta}}$ and $R\in(1,\infty)$. For $k\geq N_{1}$, we first consider the case that $h(\Delta_{k})\leq Rh(\Delta_{k-1})$. From the above inequality,
$1$ $\leq Z_{1}(\Delta_{k-1}-\Delta_{k})\Delta_{k}^{-\frac{\theta}{1-\theta}}$
$=Z_{1}(\Delta_{k-1}-\Delta_{k})h(\Delta_{k})$
$\leq RZ_{1}(\Delta_{k-1}-\Delta_{k})h(\Delta_{k-1})$
$\leq RZ_{1}\int_{\Delta_{k}}^{\Delta_{k-1}}h(s)\mathrm{d}s$
$\leq RZ_{1}\frac{1-\theta}{1-2\theta}(\Delta_{k-1}^{\frac{1-2\theta}{1-\theta}}-\Delta_{k}^{\frac{1-2\theta}{1-\theta}}).$
By defining $\mu=\frac{2\theta-1}{(1-\theta)RZ_{1}}>0$ and $\nu=\frac{1-2\theta}{1-\theta}<0$, we obtain
$0<\mu\leq\Delta_{k}^{\nu}-\Delta_{k-1}^{\nu}.$
Next, we consider the case $h(\Delta_{k})>Rh(\Delta_{k-1})$. Let $q=(\frac{1}{R})^{\frac{1-\theta}{\theta}}\in(0,1)$. From $h(\Delta_{k})=\Delta_{k}^{-\frac{\theta}{1-\theta}}>R\Delta_{k-1}^{-\frac{\theta}{1-\theta}}=Rh(\Delta_{k-1})$, we have $\Delta_{k}^{\frac{\theta}{1-\theta}}\leq\frac{1}{R}\Delta_{k-1}^{\frac{\theta}{1-\theta}}$. Taking both sides to the power of $\frac{1-\theta}{\theta}<1$, we obtain
$\Delta_{k}\leq q\Delta_{k-1}.$
Since $\nu<0$,
$\Delta_{k}^{\nu}$ $\geq q^{\nu}\Delta_{k-1}^{\nu}$
$\Delta_{k}^{\nu}-\Delta_{k-1}^{\nu}$ $\geq(q^{\nu}-1)\Delta_{k-1}^{\nu}.$
Because of $q^{\nu}-1>0$ and $\Delta_{k}\to+0$ for $k\to\infty$, we see that there exists $\bar{\mu}>0$ such that $(q^{\nu}-1)\Delta_{k-1}^{\nu}>\bar{\mu}$ for $k\geq N_{1}$. Therefore,
$\Delta_{k}^{\nu}-\Delta_{k-1}^{\nu}$ $\geq\bar{\mu}.$
Define $\hat{\mu}=\min\{\mu,\bar{\mu}\}>0$. For $k\geq N_{1}$,
$\Delta_{k}^{\nu}-\Delta_{k-1}^{\nu}$ $\geq\hat{\mu}>0.$
By summing this inequality from $k=N_{1}+1$ to $k=N>N_{1}+1$,
$\Delta_{N}^{\nu}-\Delta_{N_{1}}^{\nu}$ $\geq\hat{\mu}(N-N_{1}).$
Since $\nu<0$, there exists a positive constant $a$ such that
$\Delta_{N}$ $\leq(\Delta_{N_{1}}^{\nu}+\hat{\mu}(N-N_{1}))^{1/\nu}\leq(\hat{\mu}(N-N_{1}))^{1/\nu}\leq aN^{-\frac{1-\theta}{2\theta-1}}.$
As a result, we obtain
$\|\boldsymbol{z}^{k}-\boldsymbol{z}^{*}\|_{2}\leq\Delta_{k}\leq ak^{-\frac{1-\theta}{2\theta-1}}$
for $k\geq N_{1}$.
Next, in the case of $\theta\in(0,1/2]$, we have $\frac{1-\theta}{\theta}\geq 1$. Since $\Delta_{k-1}-\Delta_{k}\to 0$, (41) indicates that there exits a constant $Z_{2}$ for sufficiently large $k$:
$\Delta_{k}\leq Z_{2}(\Delta_{k-1}-\Delta_{k})$
$\Delta_{k}\leq\frac{Z_{2}}{1+Z_{2}}\Delta_{k-1}$
As a result, for a positive constant $Z_{3}$, it holds
$\|\boldsymbol{z}^{k}-\boldsymbol{z}^{*}\|_{2}\leq\Delta_{k}\leq Z_{3}Q^{k},$
where $Q=\frac{Z_{2}}{1+Z_{2}}\in(0,1)$.
If $\theta=0$, (40) indicates
$\|\nabla_{\boldsymbol{z}}\tilde{\ell}(\boldsymbol{z})\|_{2}\geq 1/c$
for $\boldsymbol{z}\neq\boldsymbol{z}^{*}$. Then, from (19) in lemma G.1 and (30) in lemma G.3,
$\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t+1},\boldsymbol{\theta}^{t+1})$ $\leq\ell(\boldsymbol{B}^{t},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})-C_{1}\left\|\boldsymbol{z}^{t+1}-\boldsymbol{z}^{t}\right\|_{2}^{2}$
$\leq\ell(\boldsymbol{B}^{t},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})-\frac{1}{C_{2}}\|\nabla_{\boldsymbol{z}}\tilde{\ell}(\boldsymbol{z}^{t+1})\|_{2}^{2}$
$\leq\ell(\boldsymbol{B}^{t},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})-\frac{1}{cC_{2}}$
This implies that $\boldsymbol{z}^{*}$ should be achieved by finite iterations.
As a results, the worst case of $\theta\in(1/2,1]$, $\theta\in(0,1/2]$, and $\theta=0$ can be represented as shown in Theorem 2.3.
Appendix I Discussion on $L_{1,2}$ based EIN
While we employ the $L_{0,2}$ constraint, the similar sparse selection of subgraphs can also be formulated by an $L_{1,2}$ based approach. We here briefly describe an $L_{1,2}$ based formulation of EIN, denoted as EIN-$L_{1,2}$, and discuss a few practical advantages of our $L_{0,2}$ approach compared with $L_{1,2}$.
The objective function of EIN-$L_{1,2}$ is written as
$\min_{\boldsymbol{B},\boldsymbol{b},\boldsymbol{\theta}}\$$\ell(\boldsymbol{B},\boldsymbol{b},\boldsymbol{\theta})+\lambda\|\boldsymbol{B}\|_{1,2}$ (42)
where $\lambda$ is the regularization coefficient, and $\|\boldsymbol{B}\|_{1,2}=\sum_{H\in{\mathcal{H}}}\|\boldsymbol{\beta}_{H}\|_{2}$. The similar alternating update algorithm can also be constructed for this problem. A natural counterpart of the HT update of $\boldsymbol{B}$ (3) can be defined by a proximal update
$\boldsymbol{\beta}^{t+1}_{H}$ $=\mathrm{prox}\left(\boldsymbol{\beta}_{H}-\eta\ \nabla_{\boldsymbol{\beta}_{H}}\ell^{t}\right)\text{ for }H\in{\mathcal{H}},$
where
[…]
It can be seen that, for $H\in\{H\mid\boldsymbol{\beta}^{t}_{H}=\boldsymbol{0}\}$,
$\mathrm{prox}\left(\boldsymbol{\beta}^{t}_{H}-\eta\ \nabla_{\boldsymbol{\beta}_{H}}\ell^{t}\right)=\boldsymbol{0}$
if $\|\nabla_{\boldsymbol{\beta}_{H}}\ell^{t}\|_{2}\leq\lambda$. By using the same upper bound (4), a similar pruning strategy can be defined. Assuming $H\sqsubseteq H^{\prime}$ and $H^{\prime}\in\{H\mid\boldsymbol{\beta}^{t}_{H}=\boldsymbol{0}\}$, we have
$\mathrm{UB}(H)\leq\lambda\ \Rightarrow\ \boldsymbol{\beta}^{t}_{H^{\prime}}=\boldsymbol{0}.$
In Algorithm 1, we gradually increase $s$. Instead, in EIN-$L_{1,2}$, $\lambda$ is gradually decreased, by which the strength of the sparse regularization can be gradually decreased. Since the regularization parameter is often controlled in the log space, for example, the $k$-th candidate $\lambda_{k}$ is created by $\log(\lambda_{k})=\log(\lambda_{k-1})-\Delta\lambda$, where $\lambda_{0}=\lambda_{\max}$ is the pre-specified largest value of $\lambda$, and $\Delta\lambda$ is also a pre-specified value for the decrease of $\lambda$.
Unlike the $L_{0,2}$ constraint, we cannot explicitly specify the number of non-zero components in $L_{1,2}$. This means that we cannot control the number of selected subgraphs. In the $L_{1,2}$ formulation, we observed that the number of selected subgraphs sometimes suddenly increases during optimization and when $\lambda$ changes (e.g., more than 10,000 selected subgraphs increase in the middle of optimization even when only a few subgraphs are selected in the initial several steps). A possible remedy is to control $\lambda$ adaptively to the number of selected subgraphs (i.e., if the number of selected subgraphs largely increases, decrease $\Delta\lambda$ when $k>0$ or decrease $\lambda_{0}$ when $k=0$). However, this is no more than sensible heuristics, which does not guarantee that reasonable amount of new selected subgraphs can be always generated by this approach. On the other hand, the $L_{0,2}$ approach is free from this problem, because the number of selected subgraphs can be explicitly controlled.
We empirically see that the prediction accuracy of EIN-$L_{0,2}$ and EIN-$L_{1,2}$ is not largely different. Table 5 shows a comparison based on the same setting as our main comparison Table 1. Note that for $L_{1,2}$, we used the adaptive $\Delta\lambda$ strategy (details are below).
Scheduling of $\lambda$ in $L_{1,2}$ setting
We determined $\lambda_{\max}$ by
$\lambda_{\max}=\max_{H\in{\mathcal{H}}}\|\nabla_{\boldsymbol{\beta}_{H}}\ell^{0}\|_{2}$, in which $\boldsymbol{B}^{0}$ and $\boldsymbol{b}^{0}$ were initialized by a linear model (Nakagawa et al.,, 2016) and $\boldsymbol{\theta}^{0}$ was randomly initialized. The initial value of $\Delta\lambda$ is $(\log(\lambda_{\max})-\log(0.01\lambda_{\max}))/5$. If the number of non-zero $\|\boldsymbol{\beta}_{H}\|$ increase $\geq 10$, then we update $\Delta\lambda\leftarrow 0.5\Delta\lambda$.
Table 5: Accuracy comparison between $L_{0,2}$ and $L_{1,2}$.
BZR COX2 DHFR ENZYMES PTC_MR ToxCast Cycle Cycle XOR
[…]